### v2

# Generalization Error of the Tilted Empirical Risk

Abstract
The generalization error (risk) of a supervised statistical learning algorithm quantifies its prediction ability on previously unseen data. Inspired by exponential tilting, Li et al. (2021) proposed the tilted empirical risk as a non-linear risk metric for machine learning applications such as classification and regression problems. In this work, we examine the generalization error of the tilted empirical risk. In particular, we provide uniform and information-theoretic bounds on the tilted generalization error, defined as the difference between the population risk and the tilted empirical risk, with a convergence rate of $O(1/\sqrt{n})$ where $n$ is the number of training samples. Furthermore, we study the solution to the KL-regularized expected tilted empirical risk minimization problem and derive an upper bound on the expected tilted generalization error with a convergence rate of $O(1/n)$.
[…]
1 Introduction
Empirical risk minimization (ERM) is a popular framework in machine learning. The performance of the empirical risk (ER) is affected when the data set is strongly imbalanced or contains outliers. For these scenarios, inspired by the log-sum-exponential operator with applications in multinomial linear regression and naive Bayes classifiers (Calafiore et al., 2019; Murphy, 2012; Williams and Barber, 1998), the tilted empirical risk (TER) is proposed by Li et al. (2021) for supervised learning application, such as classification and regression problems. Li et al. (2021, 2023a) showed that tilted empirical risk minimization (TERM) can handle class imbalance, mitigate the effect of outliers, and enable fairness between subgroups. Different applications of TERM have been explored, e.g., differential privacy (Lowy and Razaviyayn, 2021), semantic segmentation (Szabó et al., 2021) and noisy label self-correction (Zhou et al., 2020). In this paper, we aim to corroborate the empirical success of the TERM framework through statistical learning theory.
A central concern in statistical learning theory is understanding the efficacy of a learning algorithm when applied to test data. This evaluation is typically carried out by investigating the generalization error, which quantifies the disparity between the performance of the algorithm on the training dataset and its performance on previously unseen data, drawn from the same underlying distribution, via a risk function. Understanding the generalization behaviour of learning algorithms is one of the most important objectives in statistical learning theory. Various approaches have been developed (Rodrigues and Eldar, 2021), including VC dimension-based bounds (Vapnik, 1999), stability-based bounds (Bousquet and Elisseeff, 2002b), PAC Bayesian bounds (McAllester, 2003), and information-theoretic bounds (Russo and Zou, 2019; Xu and Raginsky, 2017). However, the TER generalization performance under unbounded loss function has not yet been studied. This paper focuses on the generalization error of the tilted empirical risk (tilted generalization error) of learning algorithms. Our contributions here can be summarized as follows,
  • We provide upper and lower bounds on the tilted generalization error via uniform and information-theoretic approaches under bounded loss functions for both positive and negative tilts and show that the convergence rates of the upper bounds are $O(1/\sqrt{n})$, as expected, where $n$ is the number of training samples.
  • We provide upper and lower bounds on the tilted generalization error under unbounded loss functions for the negative tilt with the convergence rate of $O(1/\sqrt{n})$ via a uniform approach.
  • We study the robustness of the tilted empirical risk under distribution shift induced by noise or outliers for unbounded loss functions with bounded second-moment assumption and negative tilt, and derive generalization bounds that justify the robustness properties of TERM (Li et al., 2021) for negative tilt.
  • We study the KL-regularized TERM problem and provide an upper bound on the expected tilted generalization error with convergence rate $O(1/n)$.
The paper is organised as follows: Section 2 introduces notation, the problem, and the risk functions used in this paper. Our upper and lower bounds on the tilted generalization error for bounded loss functions via uniform and information-theoretic approaches are given in Section 3. The upper and lower bounds on the tilted generalization error for unbounded loss functions with bounded second moments are given in Section 4. Section 5 is devoted to the study of robustness to the distributional shift in training samples. The KL-regularized TERM is considered in Section 6. Section 7 surveys related work. Section 8 concludes the paper. Technical tools and proofs are deferred to the appendices.
2 Preliminaries
Notations:
Upper-case letters denote random variables (e.g., $Z$), lower-case letters denote the realizations of random variables (e.g., $z$), and calligraphic letters denote sets (e.g., $\mathcal{Z}$). All logarithms are in the natural base. The tilted expectation of a random variable $X$ with tilting $\gamma$ is defined as $\frac{1}{\gamma}\log(\mathbb{E}[\exp(\gamma X)])$. The set of probability distributions (measures) over a space $\mathcal{X}$ with finite variance is denoted by $\mathcal{P}(\mathcal{X})$.
Information measures:
[…]
2.1 Problem Formulation
Let $S=\{Z_{i}\}_{i=1}^{n}$ be the training set, where each sample $Z_{i}=(X_{i},Y_{i})$ belongs to the instance space $\mathcal{Z}:=\mathcal{X}\times\mathcal{Y}$; here $\mathcal{X}$ is the input (feature) space and $\mathcal{Y}$ is the output (label) space. We assume that $Z_{i}$ are i.i.d. generated from the same data-generating distribution $\mu$.
Here we consider the set of hypothesis $\mathcal{H}$ with elements $h:\mathcal{X}\mapsto\mathcal{Y}\in\mathcal{H}$. When $\mathcal{H}$ is finite, then its cardinality is denoted by $\mathrm{card}(\mathcal{H})$. In order to measure the performance of the hypothesis $h$, we consider a non-negative loss function :×$\ell:\mathcal{H}\times\mathcal{Z}\to\mathbb{R}_{0}^{+}$.
We apply different methods to study the performance of our algorithms, including uniform and information-theoretic approaches. In uniform approaches, such as the VC-dimension and the Rademacher complexity approach (Vapnik, 1999; Bartlett and Mendelson, 2002), the hypothesis space is independent of the learning algorithm. Therefore, these methods are algorithm-independent; our results for these methods do not specify the learning algorithms.
Learning Algorithms: For information-theoretic approaches in supervised learning, following Xu and Raginsky (2017), we consider learning algorithms that are characterized by a Markov kernel (a conditional distribution) $P_{H|S}$. Such a learning algorithm maps a data set $S$ to a hypothesis in $\mathcal{H}$, which is chosen according to $P_{H|S}$. This concept thus includes randomized learning algorithms.
Robustness: Suppose that due to an outlier or noise in training samples, the underlying distribution of the training dataset, $\hat{S}$ is shifted via the distribution of noise or outlier in the training dataset, denoted as $\tilde{\mu}\in\mathcal{P}(\mathcal{Z})$. We model the distributional shift via distribution $\tilde{\mu}$ due to inspiration by the notion of influence function (Marceau and Rioux, 2001; Christmann and Steinwart, 2004; Ronchetti and Huber, 2009).
2.2 Risk Functions
The main quantity we are interested in is the population risk, defined by
$\mathrm{R}(h,\mu):=\mathbb{E}_{\tilde{Z}\sim\mu}[\ell(h,\tilde{Z})],\quad h\in\mathcal{H}.$
As the distribution $\mu$ is unknown, in classical statistical learning, the (true) population risk for $h\in\mathcal{H}$ is estimated by the (linear) empirical risk
$\widehat{\mathrm{R}}(h,S)=\frac{1}{n}\sum_{i=1}^{n}\ell(h,Z_{i}).$ (1)
The generalization error for the linear empirical risk is given by
$\mathrm{gen}(h,S):=\mathrm{R}(h,\mu)-\widehat{\mathrm{R}}(h,S);$ (2)
this is the difference between the true risk and the linear empirical risk. The TER, as a non-linear empirical risk (Li et al., 2021) and estimator of population risk, is defined by
$\widehat{\mathrm{R}}_{\gamma}(h,S)=\frac{1}{\gamma}\log\Big{(}\frac{1}{n}\sum_{i=1}^{n}\exp\big{(}\gamma\ell(h,Z_{i})\big{)}\Big{)}.$
The TER is an increasing function in $\gamma$ (Li et al., 2023a, Theorem 1), and as $\gamma\rightarrow 0$, the TER converges to the linear empirical risk in (1). Inspired by (Li et al., 2021), the primary objective is to optimize the population risk; the TERM is utilized in order to help the learning dynamics. Therefore, we decompose the population risk as follows:
$\mathrm{R}(h,\mu)=\underbrace{\mathrm{R}(h,\mu)-\widehat{\mathrm{R}}_{\gamma}(h,S)}_{\text{tilted generalization error}}+\underbrace{\widehat{\mathrm{R}}_{\gamma}(h,S),}_{\text{tilted empirical risk}}$ (3)
where we define the tilted generalization error as
$\mathrm{gen}_{\gamma}(h,S):=\mathrm{R}(h,\mu)-\widehat{\mathrm{R}}_{\gamma}(h,S).$ (4)
In learning theory, for uniform approaches, most works focus on bounding the linear generalization error $\mathrm{gen}(h,S)$ from (2) such that under the distribution of the dataset $S$, with probability at least $(1-\delta)$, it holds that for all $h\in\mathcal{H}$,
$\begin{split}|\mathrm{gen}(h,S)||\leq g(\delta,n),\end{split}$ (5)
where $g$ is a real function dependent on $\delta\in(0,1)$ and $n$ is the number of data samples. Similarly, for the tilted generalization error from (4), we are interested in finding a bound $g_{t}(\delta,n,\gamma)$ such that with probability at least $1-\delta$, under the distribution of $S$,
$\big{|}\mathrm{gen}_{\gamma}(h,S)\big{|}\leq g_{t}(\delta,n,\gamma),$ (6)
where $g_{t}$ is a real function. Furthermore, we denote the excess risk under the tilted emprical risk as,
$\mathfrak{E}_{\gamma}(\mu):=\mathrm{R}(h_{\gamma}^{*}(S),\mu)-\mathrm{R}(h^{*}(\mu),\mu),$ (7)
where $h^{*}(\mu):=\arg\min_{h\in\mathcal{H}}\mathrm{R}(h,\mu)$ and $h_{\gamma}^{*}(S):=\arg\min_{h\in\mathcal{H}}\hat{\mathrm{R}}_{\gamma}(h,S)$.
We denote the expected TER with respect to the distribution of $S$ by
$\begin{split}\overline{\mathrm{R}}_{\gamma}(h,P_{S})=\mathbb{E}_{P_{S}}[\widehat{\mathrm{R}}_{\gamma}(h,S)],\end{split}$ (8)
and we denote the tilted (true) population risk by
$\begin{split}\mathrm{R}_{\gamma}(h,P_{S})=\frac{1}{\gamma}\log\Big{(}\mathbb{E}_{P_{S}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(h,Z_{i}))\Big{]}\Big{)}.\end{split}$ (9)
Under the i.i.d. assumption, the tilted population risk is equal to an entropic risk function (Howard and Matheson, 1972). We also introduce the non-linear generalization error, which plays an important role in obtaining our bounds, as
$\widehat{\mathrm{gen}}_{\gamma}(h,S):=\mathrm{R}_{\gamma}(h,\mu)-\widehat{\mathrm{R}}_{\gamma}(h,S).$ (10)
Information-theoretic Approach: For the information-theoretic approach, as the hypothesis $H$ is a random variable under a learning algorithm as Markov kernel, i.e., $P_{H|S}$, we take expectations over the hypothesis $H$ over the above expressions for fixed $h$ to define the expected true risk, expected empirical risk, and expected tilted generalization error,
$\begin{split}\mathrm{R}(H,P_{H}\otimes\mu)&:=\mathbb{E}_{P_{H}\otimes\mu}[\ell(H,Z)],\\ \overline{\mathrm{R}}_{\gamma}(H,P_{H,S})&:=\mathbb{E}_{P_{H,S}}[\widehat{\mathrm{R}}_{\gamma}(H,S)],\\ \overline{\mathrm{gen}}_{\gamma}(H,S)&:=\mathbb{E}_{P_{H,S}}[\mathrm{gen}_{\gamma}(H,S)].\end{split}$ (11)
Similar to (11), we define the tilted population risk and the tilted generalization error, (9) and (10), as expectations. We also provide upper bounds on the absolute value of expected tilted generalization error with respect to the joint distribution of $S$ and $H$, of the form
$|\overline{\mathrm{gen}}_{\gamma}(H,S)|\leq g_{e}(n,\gamma),$
where $g_{e}$ is a real function.
3 Generalization Bounds for Bounded Loss Function
Upper bounds under linear empirical risk for bounded loss functions via information theoretic and uniform bounds are studied by Shalev-Shwartz and Ben-David (2014) and Xu and Mannor (2012), respectively. Inspired by these works, in this section, we provide upper bounds on the tilted generalization error via uniform and information-theoretic approaches for bounded loss functions with the convergence rate of $O(1/\sqrt{n})$ which is similar to generalization error under linear empirical risk. Upper bounds via stability (Bousquet and Elisseeff, 2002b), Rademacher complexity (Bartlett and Mendelson, 2002) and PAC-Bayesian approaches (Alquier, 2021) are provided in Appendix H. All proof details are deferred to Appendix D.
In this section the following assumption is made.
Assumption 3.1 (Bounded loss function).
There is a constant $M$ such that the loss function, $(h,z)\mapsto\ell(h,z)$ satisfies $0\leq\ell(h,z)\leq M$ uniformly for all $h\in\mathcal{H},z\in\mathcal{Z}$.
Assumption 3.1 will be relaxed in Section 4.
3.1 Uniform Bounds
For uniform bounds of the type (6) we decompose the tilted generalization error (4) as follows,
$\begin{split}&\mathrm{gen}_{\gamma}(h,S)=\underbrace{\mathrm{R}(h,\mu)-\mathrm{R}_{\gamma}(h,\mu^{\otimes n})}_{I_{1}}\\ &\qquad+\underbrace{\mathrm{R}_{\gamma}(h,\mu^{\otimes n})-\widehat{\mathrm{R}}_{\gamma}(h,S)}_{I_{2}},\end{split}$ (12)
where $I_{1}$ is the difference between the population risk and the tilted population risk and $I_{2}$ is the non-linear generalization error.
We first derive an upper bound on term $I_{1}$ in the following Proposition.
Proposition 3.2.
Under Assumption 3.1, for $\gamma\in\mathbb{R}$ the difference between the population risk and the tilted population risk satisfies
$\begin{split}&\frac{-1}{2\gamma}\mathrm{Var}\big{(}\exp(\gamma\ell(h,Z))\big{)}\leq\mathrm{R}(h,\mu)-\mathrm{R}_{\gamma}(h,\mu^{\otimes n})\\ &\leq\frac{-\exp(-2\gamma M)}{2\gamma}\mathrm{Var}\big{(}\exp(\gamma\ell(h,Z))\big{)}.\end{split}$ (13)
Note that for $\gamma\rightarrow 0$, the upper and lower bounds in Proposition 3.2 are zero.
As the log function is Lipschitz on a bounded interval, applying the Hoeffding inequality to term $I_{2}$ and Proposition 3.2 to term $I_{1}$ in (12), we obtain the following upper bound on the tilted generalization error.
Theorem 3.3.
Given Assumption 3.1, for any fixed $h\in\mathcal{H}$ with probability at least $(1-\delta)$ the tilted generalization error satisfies the upper bound,
$\begin{split}\mathrm{gen}_{\gamma}(h,S)&\leq\frac{-\exp(-2\gamma M)}{2\gamma}\operatorname{Var}\big{(}\exp(\gamma\ell(h,Z))\big{)}\\ &\qquad+\frac{\big{(}\exp(|\gamma|M)-1\big{)}}{|\gamma|}\sqrt{\frac{\log(2/\delta)}{2n}}.\end{split}$
Theorem 3.4.
Under the same assumptions of Theorem 3.3, for a fixed $h\in\mathcal{H}$, with probability at least $(1-\delta)$, the tilted generalization error satisfies the lower bound
$\begin{split}\mathrm{gen}_{\gamma}(h,S)&\geq\frac{-1}{2\gamma}\operatorname{Var}\big{(}\exp(\gamma\ell(h,Z))\big{)}\\ &\qquad-\frac{\big{(}\exp(|\gamma|M)-1\big{)}}{|\gamma|}\sqrt{\frac{\log(2/\delta)}{2n}}.\end{split}$ (14)
Remark 3.5.
The lower bound in Theorem 3.4 for a negative $\gamma$ can be tighter than the lower bound on the tilted generalization error for a positive $\gamma$, (14), as the variance is positive.
Combining Theorem 3.3 and Theorem 3.4, we derive an upper bound on the absolute value of the titled generalization error.
Corollary 3.6.
Let $A(\gamma)=(1-\exp(\gamma M))^{2}$. Under the same assumptions in Theorem 3.3, with probability at least $(1-\delta)$, and a finite hypothesis space, the absolute value of the titled generalization error satisfies
$\begin{split}&\sup_{h\in\mathcal{H}}|\mathrm{gen}_{\gamma}(h,S)|\\ &\leq\frac{\big{(}\exp(|\gamma|M)-1\big{)}}{|\gamma|}\sqrt{\frac{\log(\mathrm{card}(\mathcal{H}))+\log(2/\delta)}{2n}}\\ &\qquad+\frac{\max(1,\exp(-2\gamma M))A(\gamma)}{8|\gamma|},\end{split}$
where $A(\gamma)=(1-\exp(\gamma M))^{2}$.
Corollary 3.7.
Under the assumptions in Corollary 3.6, if $\gamma$ is of order $O(n^{-\beta})$ for $\beta>0$, then, as $A(\gamma)\sim\gamma^{2}M^{2}$ for $\gamma\rightarrow 0$, the upper bound on the absolute tilted generalization error in Corollary 3.6 has a convergence rate of $\max\big{(}O(1/\sqrt{n}),O(n^{-\beta})\big{)}$ as $n\rightarrow\infty$.
Remark 3.8.
Choosing $\beta\geq 1/2$ in Corollary 3.7 gives a convergence rate of $O(1/\sqrt{n})$ for the tilted generalization error.
Remark 3.9 (The influence of $\gamma$).
As $\gamma\rightarrow 0$, the upper bound in Corollary 3.6 on the absolute value of tilted generalization error converges to the upper bound on absolute value of the generalization error under the ERM algorithm obtained by Shalev-Shwartz and Ben-David (2014),
$\sup_{h\in\mathcal{H}}|\mathrm{gen}(h,S)|\leq M\sqrt{\frac{\log(\mathrm{card}(\mathcal{H}))+\log(2/\delta)}{2n}}.$ (15)
In particular, ${\big{(}\exp(|\gamma|M)-1\big{)}}/{|\gamma|}\rightarrow M$ and the first term in Corollary 3.6 vanishes. Therefore, the upper bound converges to a uniform bound on the linear empirical risk.
Remark 3.10.
The term $-\exp(-2\gamma M)/\gamma$ in Theorem 3.3 can cause the upper bounds in Theorem 3.3 to be tighter for $\gamma>0$ than for $\gamma<0$. Furthermore, in comparison with the uniform upper bound on the linear generalization error, (15), we obtain a tilted generalization error upper bounds which can be tighter, for sufficiently large values of samples $n$ and small values of $\gamma$.
Using Corollary 3.7, we derive an upper bound on the excess risk.
Corollary 3.11.
Under the same assumptions in Theorem 3.3, and a finite hypothesis space, with probability at least $(1-\delta)$, the excess risk of tilted empirical risk satisfies
$\begin{split}\mathfrak{E}_{\gamma}(\mu)&\leq\frac{2\big{(}\exp(|\gamma|M)-1\big{)}}{|\gamma|}\sqrt{\frac{\log(\mathrm{card}(\mathcal{H}))+\log(2/\delta)}{2n}}\\ &\qquad+\frac{2\max(1,\exp(-2\gamma M))A(\gamma)}{8|\gamma|},\end{split}$
where $A(\gamma)=(1-\exp(\gamma M))^{2}$.
The theorems in this section assumed that the hypothesis space is finite; this is for example the case in classification problems with a finite number of classes. If this assumption is violated, we can apply the growth function technique from (Bousquet et al., 2003; Vapnik, 1999). Furthermore, the growth function can be bounded by VC-dimension in binary classification (Vapnik, 1999) or Natarajan dimension (Holden and Niranjan, 1995) for multi-class classification scenarios. Note that the VC-dimension and Rademacher complexity bounds are uniform bounds and are independent of the learning algorithms.
3.2 Information-theoretic Bounds
Next, we provide an upper bound on the expected tilted generalization error. All proof details are deferred to Appendix D.2. For information-theoretic bounds, we employ the following decomposition of the expected tilted generalization error,
$\begin{split}&\overline{\mathrm{gen}}_{\gamma}(H,S)\\ &=\{\mathrm{R}(H,P_{H}\otimes\mu)-\overline{\mathrm{R}}_{\gamma}(H,P_{H}\otimes\mu^{\otimes n})\}\\ &+\{\overline{\mathrm{R}}_{\gamma}(H,P_{H}\otimes\mu^{\otimes n})-\overline{\mathrm{R}}_{\gamma}(H,P_{H,S})\}.\end{split}$ (16)
The following is helpful in deriving the upper bound.
Proposition 3.12.
Under Assumption 3.1, the following inequality holds for any learning algorithm, $P_{H|S}$,
$\begin{split}&\Big{|}\overline{\mathrm{R}}_{\gamma}(H,P_{H}\otimes\mu^{\otimes n})-\overline{\mathrm{R}}_{\gamma}(H,P_{H,S})\Big{|}\\ &\leq\frac{(\exp(|\gamma|M)-1)}{|\gamma|}\sqrt{\frac{I(H;S)}{2n}}.\end{split}$ (17)
Using Proposition 3.12, we derive the following upper and lower bounds on the expected generalization error.
Theorem 3.13.
Under Assumption 3.1, the expected tilted generalization error satisfies
$\overline{\mathrm{gen}}_{\gamma}(H,S)\leq\frac{(\exp(|\gamma|M)-1)}{|\gamma|}\sqrt{\frac{I(H;S)}{2n}}-\frac{\gamma\exp(-\gamma M)}{2}\Big{(}1-\frac{1}{n}\Big{)}\mathbb{E}_{P_{H}}\big{[}\mathrm{Var}_{\tilde{Z}\sim\mu}(\ell(H,\tilde{Z}))\big{]}.$ (18)
Theorem 3.14.
Under the same assumptions in Theorem 3.13, the expected tilted generalization error satisfies
$\begin{split}&\overline{\mathrm{gen}}_{\gamma}(H,S)\geq-\frac{(\exp(|\gamma|M)-1)}{|\gamma|}\sqrt{\frac{I(H;S)}{2n}}\\ &\qquad-\frac{\gamma\exp(\gamma M)}{2}\Big{(}1-\frac{1}{n}\Big{)}\mathbb{E}_{P_{H}}\big{[}\mathrm{Var}_{\tilde{Z}\sim\mu}(\ell(H,\tilde{Z}))\big{]}.\end{split}$
Combining Theorem 3.13 and Theorem 3.14, we derive an upper bound on the absolute value of the expected tilted generalization error.
Corollary 3.15.
Under the same assumptions in Theorem 3.13, the absolute value of the expected titled generalization error satisfies
$\begin{split}|\overline{\mathrm{gen}}_{\gamma}(H,S)|&\leq\frac{(\exp(|\gamma|M)-1)}{|\gamma|}\sqrt{\frac{I(H;S)}{2n}}\\ &\qquad+\frac{|\gamma|M^{2}\exp(|\gamma|M)}{8}\Big{(}1-\frac{1}{n}\Big{)}.\end{split}$
Remark 3.16.
In Corollary 3.15, we observe that by choosing $\gamma=O(n^{-\beta})$, the overall convergence rate of the generalization error upper bound is $\max(O(1/\sqrt{n}),O(n^{-\beta}))$ for bounded $I(H;S)$. For $\beta\geq 1/2$, the convergence rate of (18) is the same as the convergence rate of the expected upper bound in (Xu and Raginsky, 2017). In addition, for $\gamma\rightarrow 0$, the upper bound in Corollary 3.15 converges to the expected upper bound in (Xu and Raginsky, 2017).
The results in this section are non-vacuous for bounded $I(H;S)$. If this assumption is violated, we can apply the individual sample method (Bu et al., 2020), chaining methods (Asadi et al., 2018), or conditional mutual information frameworks (Steinke and Zakynthinou, 2020) to derive tighter upper bound for the tilted generalization error.
4 Generalization Bounds for Unbounded Loss Functions
In the previous section, we assumed that the loss function is bounded which is limiting in practice. Several works already proposed some solutions to overcome the boundedness assumption under linear empirical risk (Lugosi and Neu, 2022; Haddouche and Guedj, 2022; Alquier and Guedj, 2018; Holland, 2019) via a PAC-Bayesian approach. In this section, we derive upper bounds on the tilted generalization error via uniform approach for negative tilt ($\gamma<0$) under bounded second moment assumption with convergence rate of $O(1/\sqrt{n})$. In particular, we relax the bounded assumption (Assumption 3.1). The following assumptions are made for the uniform analysis.
Assumption 4.1 (Uniform bounded second moment).
There is a constant $\kappa_{u}\in\mathbb{R}^{+}$ such that the loss function $(H,Z)\mapsto\ell(H,Z)$ satisfies $\mathbb{E}_{\mu}[\ell^{2}(h,Z)]\leq\kappa_{u}^{2}$ uniformly for all $h\in\mathcal{H}$.
The assumption on second moment, Assumption 4.1, can be satisfied if the loss function is sub-Gaussian or sub-Exponential (Boucheron et al., 2013) under the distribution $\mu$ for all $h\in\mathcal{H}$. All proof details for the results in this section are deferred to Appendix E. In this following we present the results based on uniform bounds. The results for unbounded loss function based on information theoretic bounds are presented in Appendix E.2.
For unbounded loss function, we consider the decomposition of the tilted generalization error in (12) for $\gamma<0$. The term $I_{1}$ can be bounded using Lemma C.9. Then, we apply Bernstein’s inequality (Boucheron et al., 2013) to provide upper and lower bounds on the second term, $I_{2}$.
Theorem 4.2.
Given Assumption 4.1, for any fixed $h\in\mathcal{H}$ with probability at least $(1-\delta)$, then the following upper bound holds on the tilted generalization error for $\gamma<0$,
$\begin{split}\mathrm{gen}_{\gamma}(h,S)&\leq 2\kappa_{u}\exp(-\gamma\kappa_{u})\sqrt{\frac{\log(2/\delta)}{n}}\\ &-\frac{4\exp(-\gamma\kappa_{u})\log(2/\delta)}{3n\gamma}-\frac{\gamma}{2}\kappa_{u}^{2}.\end{split}$ (19)
Theorem 4.3.
Given Assumption 4.1, there exists a $\zeta\in(0,1)$ such that for $n\geq\frac{(4\gamma^{2}\kappa_{u}^{2}+8/3\zeta)\log(2/\delta)}{\zeta^{2}\exp(2\gamma\kappa_{u})}$, for any fixed $h\in\mathcal{H}$ with probability at least $(1-\delta)$, and $\gamma<0$, the following lower bound on the tilted generalization error holds
$\begin{split}\mathrm{gen}_{\gamma}(h,S)&\geq-\frac{2\kappa_{u}\exp(-\gamma\kappa_{u})}{(1-\zeta)}\sqrt{\frac{\log(2/\delta)}{n}}\\ &+\frac{4\exp(-\gamma\kappa_{u})(\log(2/\delta))}{3n\gamma(1-\zeta)}.\end{split}$ (20)
Combining Theorem 4.2 and Theorem 4.3, we derive an upper bound on the absolute value of the tilted generalization error.
Corollary 4.4.
Under the same assumptions in Theorem 4.3 and a finite hypothesis space, then for $n\geq\frac{(4\gamma^{2}\kappa_{u}^{2}+8/3\zeta)\log(2/\delta)}{\zeta^{2}\exp(2\gamma\kappa_{u})}$, for $\gamma<0$ and with probability at least $(1-\delta)$, the absolute value of the titled generalization error satisfies
$\sup_{h\in\mathcal{H}}|\mathrm{gen}_{\gamma}(h,S)|\leq\frac{2\kappa_{u}\exp(-\gamma\kappa_{u})}{(1-\zeta)}\sqrt{\frac{B(\delta)}{n}}-\frac{4\exp(-\gamma\kappa_{u})B(\delta)}{3n\gamma(1-\zeta)}-\frac{\gamma}{2}\kappa_{u}^{2},$ (21)
where $B(\delta)=\log(\mathrm{card}(\mathcal{H}))+\log(2/\delta)$.
Remark 4.5.
For $\gamma\asymp n^{-1/2}$, $n\geq\frac{(4\gamma^{2}\kappa_{u}^{2}+8/3\zeta)\log(2/\delta)}{\zeta^{2}\exp(-2\kappa_{u})}$ and $-1<\gamma<0$, the upper bound in Corollary 4.4 gives a theoretical guarantee on the convergence rate of $O(n^{-1/2})$. Using TER with negative $\gamma$ can help to derive an upper bound on the absolute value of the tilted generalization error under the bounded second-moment assumption.
Similar to Corollary 3.11, an upper bound on excess risk under unbounded loss function assumption can be derived.
Corollary 4.6.
Under the same assumption in Theorem 4.3 and a finite hypothesis space, then for $n\geq\frac{(4\gamma^{2}\kappa_{u}^{2}+8/3\zeta)\log(2/\delta)}{\zeta^{2}\exp(2\gamma\kappa_{u})}$ with probability at least $(1-\delta)$ and $\gamma<0$, the excess risk of tilted empirical risk satisfies
$\mathfrak{E}_{\gamma}(\mu)$ $\leq\frac{4\kappa_{u}\exp(-\gamma\kappa_{u})}{(1-\zeta)}\sqrt{\frac{B(\delta)}{n}}-\gamma\kappa_{u}^{2}-\frac{8\exp(-\gamma\kappa_{u})B(\delta)}{3n\gamma(1-\zeta)},$ (22)
where $B(\delta)=\log(\mathrm{card}(\mathcal{H}))+\log(2/\delta)$.
5 Robustness of TERM
As shown in experiments by Li et al. (2021), the tilted empirical risk is robust to noise or outliers samples during training using negative tilt ($\gamma<0$). In this section, we study the robustness of the TER under distributional shift, $\tilde{\mu}$ and the following assumption is made.
Assumption 5.1 (Uniform bounded second moment under $\tilde{\mu}$).
There is a constant $\kappa_{s}\in\mathbb{R}^{+}$ such that the loss function $(H,Z)\mapsto\ell(H,Z)$ satisfies $\mathbb{E}_{\tilde{\mu}}[\ell^{2}(h,Z)]\leq\kappa_{s}^{2}$ uniformly for all $h\in\mathcal{H}$.
Using the functional derivative (Cardaliaguet et al., 2019), we can provide the following results.
Proposition 5.2.
Given Assumption 5.1 and Assumption 4.1, then the difference of tilted population risk under, (9), between $\mu$ and $\tilde{\mu}$ is bounded as follows,
$\frac{1}{|\gamma|}\Big{|}\log(\mathbb{E}_{\tilde{Z}\sim\mu}[C(h,\tilde{Z})])-\log(\mathbb{E}_{\tilde{Z}\sim\tilde{\mu}}[C(h,\tilde{Z})])\Big{|}\leq\frac{\mathbb{TV}(\mu,\tilde{\mu})}{\gamma^{2}}\frac{\exp(|\gamma|\kappa_{u})-\exp(|\gamma|\kappa_{s})}{(\kappa_{u}-\kappa_{s})},$ (23)
where $C(h,\tilde{Z})=\exp(\gamma\ell(h,\tilde{Z}))$.
Note that, for positive $\gamma$, the result in Proposition 5.2 does not hold and can be unbounded. Using Proposition 5.2, we can provide an upper and lower bounds on the tilted generalization error under distributional shift.
Proposition 5.3.
Given Assumptions  4.1 and 5.1, for any fixed $h\in\mathcal{H}$ and with probability least $(1-\delta)$ for $\gamma<0$, then the following upper bound holds on the tilted generalization error
$\begin{split}&\mathrm{gen}_{\gamma}(h,\hat{S})\leq 2\kappa_{s}\exp(-\gamma\kappa_{s})\sqrt{\frac{\log(2/\delta)}{n}}-\frac{\gamma}{2}\kappa_{u}^{2}\\ &\quad-\frac{4\exp(-\gamma\kappa_{s})(\log(2/\delta))}{3n\gamma}+\frac{\mathbb{TV}(\mu,\tilde{\mu})}{\gamma^{2}}D(\kappa_{s},\kappa_{u}),\end{split}$
where $D(\kappa_{s},\kappa_{u})=\frac{\exp(|\gamma|\kappa_{u})-\exp(|\gamma|\kappa_{s})}{(\kappa_{u}-\kappa_{s})}$ and $\hat{S}$ is the training dataset under the distributional shift.
Proposition 5.4.
Given Assumptions  4.1 and 5.1, for any fixed $h\in\mathcal{H}$ and with probability least $(1-\delta)$, there exists a $\zeta\in(0,1)$ such that for $n\geq\frac{(4\gamma^{2}\kappa_{u}^{2}+8/3\zeta)\log(2/\delta)}{\zeta^{2}\exp(2\gamma\kappa_{u})}$ and $\gamma<0$, such that the following upper bound holds on the tilted generalization error
$\begin{split}&\mathrm{gen}_{\gamma}(h,\hat{S})\geq-\frac{2\kappa_{s}\exp(-\gamma\kappa_{s})}{(1-\zeta)}\sqrt{\frac{\log(2/\delta)}{n}}\\ &\qquad+\frac{4\exp(-\gamma\kappa_{s})(\log(2/\delta))}{3n\gamma(1-\zeta)}-\frac{\mathbb{TV}(\mu,\tilde{\mu})}{\gamma^{2}}D(\kappa_{s},\kappa_{u}),\end{split}$
where $D(\kappa_{s},\kappa_{u})=\frac{\exp(|\gamma|\kappa_{u})-\exp(|\gamma|\kappa_{s})}{(\kappa_{u}-\kappa_{s})}$ and $\hat{S}$ is the training dataset under the distributional shift.
Combining Proposition 5.3 and Proposition 5.4, we derive an upper bound on the absolute value of the tilted generalization error under distributional shift.
Theorem 5.5.
Under the same assumptions in Proposition 5.4, then for $n\geq\frac{(4\gamma^{2}\kappa_{u}^{2}+8/3\zeta)\log(2/\delta)}{\zeta^{2}\exp(2\gamma\kappa_{u})}$ and $\gamma<0$, and with probability at least $(1-\delta)$, the absolute value of the tilted generalization error under distributional shift satisfies
$\begin{split}&\sup_{h\in\mathcal{H}}|\mathrm{gen}_{\gamma}(h,\hat{S})|\leq\frac{2\kappa_{s}\exp(-\gamma\kappa_{s})}{(1-\zeta)}\sqrt{\frac{B(\delta)}{n}}-\frac{4\exp(-\gamma\kappa_{s})B(\delta)}{3n\gamma(1-\zeta)}\\ &\qquad-\frac{\gamma\kappa_{u}^{2}}{2}+\frac{\mathbb{TV}(\mu,\tilde{\mu})}{\gamma^{2}}\frac{\big{(}\exp(|\gamma|\kappa_{u})-\exp(|\gamma|\kappa_{s})\big{)}}{(\kappa_{u}-\kappa_{s})},\end{split}$ (24)
where $B(\delta)=\log(\mathrm{card}(\mathcal{H}))+\log(2/\delta)$.
Similar results to Proposition 5.3, can be derived via an information-theoretic approach.
Robustness vs Generalization: The term $\frac{\mathbb{TV}(\mu,\tilde{\mu})}{\gamma^{2}}\frac{\big{(}\exp(|\gamma|\kappa_{u})-\exp(|\gamma|\kappa_{s})\big{)}}{(\kappa_{u}-\kappa_{s})}$ represents the distributional shift cost (or robustness) associated with the TER. This cost can be reduced by increasing $|\gamma|$. However, increasing $|\gamma|$ also amplifies other terms in the upper bound of the tilted generalization error. Therefore, there is a trade-off between robustness and generalization, particularly for $\gamma<0$ in the TER. Interestingly, Li et al. (2021) also observed this trade-off for negative tilt.
6 The KL-Regularized TERM Problem
Our upper bound in Corollary 3.15 on the absolute value of expected generalization error depends on the mutual information between $H$ and $S$. Therefore, it is of interest to investigate an algorithm which minimizes the regularized expected TERM via mutual information.
$P_{H|S}^{\star}=\mathop{\mathrm{arg}\,\inf}_{P_{H|S}}\overline{\mathrm{R}}_{\gamma}(H,P_{H,S})+\frac{1}{\alpha}I(H;S),$ (25)
where $\alpha$ is the inverse temperature. As discussed by Xu and Raginsky (2017); Aminian et al. (2023), the regularization problem in (25) is dependent on the data distribution, $P_{S}$. Therefore, we relax the problem in (25) by considering the following regularized version via KL divergence,
$P_{H|S}^{\star}=\mathop{\mathrm{arg}\,\inf}_{P_{H|S}}\overline{\mathrm{R}}_{\gamma}(H,P_{H,S})+\frac{1}{\alpha}\mathrm{KL}(P_{H|S}\|\pi_{H}|P_{S}),$ (26)
where $I(H;S)\leq\mathrm{KL}(P_{H|S}\|\pi_{H}|P_{S})$ and $\pi_{H}$ is a prior distribution over hypothesis space $\mathcal{H}$. All proof details are deferred to Appendix G.
Proposition 6.1.
The solution to the expected TERM regularized via KL divergence, (26), is the tilted Gibbs Posterior (a.k.a. Gibbs Algorithm),
$P_{H|S}^{\gamma}:=\frac{\pi_{H}}{F_{\alpha}(S)}\Big{(}\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))\Big{)}^{-\alpha/\gamma},$ (27)
where $F_{\alpha}(S)$ is a normalization factor.
Note that the Gibbs posterior,
$P_{H|S}^{\alpha}:=\frac{\pi_{H}}{\tilde{F}_{\alpha}(S)}\exp\Big{(}-\alpha\Big{(}\frac{1}{n}\sum_{i=1}^{n}\ell(H,Z_{i})\Big{)}\Big{)},$ (28)
is the solution to the KL-regularized ERM minimization problem, where $\tilde{F}_{\alpha}(S)$ is the normalization factor. Therefore, the tilted Gibbs posterior is different from the Gibbs posterior, (28). It can be shown that for $\gamma\rightarrow 0$, the tilted Gibbs posterior converges to the Gibbs posterior. Therefore, it is interesting to study the expected generalization error of the tilted Gibbs posterior. For this purpose, we give an exact characterization of the difference between the expected TER under the joint and the product of marginal distributions of $H$ and $S$. More results regarding the Gibbs posterior are provided in Appendix G.
Proposition 6.2.
The difference between the expected TER under the joint and product of marginal distributions of $H$ and $S$ can be expressed as,
$\begin{split}&\overline{\mathrm{R}}_{\gamma}(H,P_{H}\otimes\mu)-\overline{\mathrm{R}}_{\gamma}(H,P_{H,S})=\frac{I_{\mathrm{SKL}}(H;S)}{\alpha}.\end{split}$ (29)
We next provide a parametric upper bound on the tilted generalization error of the tilted Gibbs posterior.
Theorem 6.3.
Under Assumption 3.1, the expected generalization error of the tilted Gibbs posterior satisfies
$\begin{split}&\overline{\mathrm{gen}}_{\gamma}(H,S)\leq\frac{\alpha(\exp(|\gamma|M)-1)^{2}}{2\gamma^{2}n}\\ &-\frac{\gamma\exp(-\gamma M)}{2}\big{(}1-\frac{1}{n}\big{)}\mathbb{E}_{P_{H}}\big{[}\mathrm{Var}_{\tilde{Z}\sim\mu}(\ell(H,\tilde{Z}))\big{]}.\end{split}$ (30)
Similar to Corollary 3.15, we derive the following upper bound on the absolute value of the expected tilted generalization error of the tilted generalization error.
Corollary 6.4.
Under the same assumptions in Theorem 6.3, the absolute value of the expected tilted generalization error of the tilted Gibbs posterior satisfies
$\begin{split}&|\overline{\mathrm{gen}}_{\gamma}(H,S)|\leq\frac{\alpha(\exp(|\gamma|M)-1)^{2}}{2\gamma^{2}n}\\ &+\frac{|\gamma|M^{2}\exp(|\gamma|M)}{8}\Big{(}1-\frac{1}{n}\Big{)}.\end{split}$ (31)
Remark 6.5 (Convergence rate).
If $\gamma=O(1/n)$, then we obtain a theoretical bound on the convergence rate of $O(1/n)$ for the upper bound on the tilted generalization error of the tilted Gibbs posterior.
Remark 6.6 (Discussion of $\gamma$).
From the upper bound in Theorem 6.3, we can observe that under $\gamma\rightarrow 0$ and Assumption 3.1, the upper bound converges to the upper bound on the Gibbs posterior (Aminian et al., 2021a). For positive tilt $(\gamma>0)$, and sufficient large value of $n$, the upper bound in Theorem 6.3, can be tighter than the upper bound on the Gibbs posterior.
7 Related Works
Generalization Error Analysis: Different approaches have been applied to study the generalization error of general learning problems under empirical risk minimization, including VC dimension-based, Rademacher complexity, PAC-Bayesian, stability and information-theoretic bounds. In this section, we discuss the related works about Uniform and information-theoretic bounds. More related works for generalization error analysis are discussed in Appendix B.
Uniform Bounds: Uniform bounds (or VC bounds) are proposed by  Vapnik and Chervonenkis (1971); Bartlett et al. (1998, 2019). For any class of functions $\mathcal{F}$ of VC dimension $d$, with probability at least $1-\delta$ the generalization error is $O\big{(}(d+\log(1/\delta))^{1/2}n^{-1/2}\big{)}$. This bound depends solely on the VC dimension of the function class and on the sample size; in particular, it is independent of the learning algorithm.
Information-theoretic bounds: Russo and Zou (2019); Xu and Raginsky (2017) propose using the mutual information between the input training set and the output hypothesis to upper bound the expected generalization error. Multiple approaches have been proposed to tighten the mutual information-based bound: Bu et al. (2020) provide tighter bounds by considering the individual sample mutual information; Asadi et al. (2018); Asadi and Abbe (2020) propose using chaining mutual information; and Steinke and Zakynthinou (2020); Hafez-Kolahi et al. (2020); Aminian et al. (2020, 2021b) provide different upper bounds on the expected generalization error based on the linear empirical risk framework.
The aforementioned approaches are applied to study the generalization error in the linear empirical risk framework. To our knowledge, the generalization error of the tilted empirical risk minimization from information theoretical and uniform approach perspectives has not been explored.
8 Conclusion and Future Work
In this paper, we study the tilted empirical risk minimization, as proposed by Li et al. (2021). In particular, we established an upper and lower bound on the tilted generalization error of the tilted empirical risk through uniform and information-theoretic approaches, obtaining theoretical guarantees that the convergence rate is $O(1/\sqrt{n})$ under bounded loss functions for negative and positive tilts. Furthermore, we provide an upper bound on the titled generalization error for the unbounded loss function. We also study the tilted generalization error under distribution shift in the training dataset due to noise or outliers, where we discussed the generalization and robustness trade-off. Additionally, we explore the KL-regularized tilted empirical risk minimization, where the solution involves the tilted Gibbs posterior, and we derive a parametric upper bound on this minimization with a convergence rate of $O(1/n)$ under some conditions.
Our current results are applicable to scenarios with bounded loss functions and in scenarios with negative tilting ($\gamma<0$) when the possibly unbounded loss function has a bounded second moment. However, our current results for the asymptotic regime $\gamma\rightarrow\infty$, where the tilted empirical risk is equal to maximum loss, are vacuous. Therefore, studying the generalization performance of tilted empirical risk minimization under unbounded loss functions for positive tilting, and obtaining informative bounds in the asymptotic regime $\gamma\rightarrow\infty$, are planned as future work. Moreover, we intend to apply other approaches from the literature, such as those discussed by Aminian et al. (2023), to derive further bounds on the tilted generalization error in the mean-field regime.
Acknowledgements
Gholamali Aminian, Gesine Reinert and Samuel N. Cohen acknowledge the support of the UKRI Prosperity Partnership Scheme (FAIR) under EPSRC Grant EP/V056883/1 and the Alan Turing Institute. Gesine Reinert is also supported in part by EPSRC grants EP/W037211/1 and EP/R018472/1. Samuel N. Cohen also acknowledges the support of the Oxford–Man Institute for Quantitative Finance. Amir R. Asadi is supported by Leverhulme Trust grant ECF-2023-189 and Isaac Newton Trust grant 23.08(b).
Appendix A Overview of Main Results
An overview of our main results is provided in Fig. 1. All notations are summarized in Table 1.
Table 1: Summary of notations in the paper
Notation Definition Notation Definition
$S$ Training set $Z_{i}$ $i$-th sample
$\mathcal{X}$ Input (feature) space $\mathcal{Y}$ Output (label) space
$\mu$ Data-generating distribution $\tilde{\mu}$ Data-generating distribution under distributional shift
$\gamma$ Tilt parameter $\mathrm{KL}(P\|Q)$ KL-divergence between $P$ and $Q$
$I(X;Y)$ Mutual information $D_{\mathrm{SKL}}(P\|Q)$ Symmetrized KL divergence
$\mathbb{TV}(P,Q)$ Total variation distance $\kappa_{u}$ Bound on second moment
$h$ Hypothesis $\mathcal{H}$ Hypothesis space
$\ell(h,z)$ Loss function $R(h,\mu)$ Population (true) risk
$\widehat{\mathrm{R}}_{\gamma}(h,S)$ Tilted empirical risk $\mathrm{gen}_{\gamma}(h,S)$ Tilted generalization error
$\overline{\mathrm{R}}_{\gamma}(h,P_{S})$ Tilted population risk $\mathfrak{E}_{\gamma}(\mu)$ Excess risk
Tilted Generalization Error Bounded Loss Function ($\gamma\in\mathbb{R}$) Section 3 Information-theoretic Bound Upper bound (Theorem 3.13) & Lower bound (Theorem 3.14) Uniform Bound Upper bound (Theorem 3.3) & Lower bound (Theorem 3.4) Unbounded Loss Function ($\gamma\in\mathbb{R}_{<0}$) Section 4 Information-theoretic Bound Upper bound (Theorem E.4) & Lower bound (Theorem E.6) Uniform Bound Upper bound (Theorem 4.2) & Lower bound (Theorem 4.3) Robustness of TERM ($\gamma\in\mathbb{R}_{<0}$) Section 5 Uniform Bound Upper bound (Proposition 5.3) & Lower bound (Theorem 5.4) KL-Regularized TERM ($\gamma\in\mathbb{R}$) Section 6 Information-theoretic Bound Upper bound (Theorem 6.3)
Figure 1: Overview of the main results
Appendix B Other Related Works
This section details related works about Tilted empirical risk minimization, the Rademacher complexity, stability and PAC-Bayesian bounds, as well as related work on unbounded loss functions.
Rademacher Complexity Bounds: This approach is a data-dependent method to provide an upper bound on the generalization error based on the Rademacher complexity of the function class $\mathcal{H}$, see (Bartlett and Mendelson, 2002; Golowich et al., 2018). Bounding the Rademacher complexity involves the model parameters. Typically, in Rademacher complexity analysis, a symmetrization technique is used which can be applied to the empirical risk, but not directly to the TER.
Stability Bounds: Stability-based bounds for generalization error are given in Bousquet and Elisseeff (2002a); Bousquet et al. (2020); Mou et al. (2017); Chen et al. (2018); Aminian et al. (2023). For stability analysis, the key tool is Lemma 7 in Bousquet and Elisseeff (2002b), which is based on ERM linearity. Therefore, we can not apply stability analysis to TER directly.
PAC-Bayes bounds: First proposed by Shawe-Taylor and Williamson (1997); McAllester (1999) and McAllester (2003), PAC-Bayesian analysis provides high probability bounds on the generalization error in terms of the KL divergence between the data-dependent posterior induced by the learning algorithm and a data-free prior that can be chosen arbitrarily (Alquier, 2021). There are multiple ways to generalize the standard PAC-Bayesian bounds, including using information measures other than KL divergence (Alquier and Guedj, 2018; Bégin et al., 2016; Hellström and Durisi, 2020; Aminian et al., 2021b) and considering data-dependent priors (Rivasplata et al., 2020; Catoni, 2007; Dziugaite and Roy, 2018; Ambroladze et al., 2007). However, this method has not been applied to TER to provide generalization error bounds.
Tilted Empirical Risk Minimization: The TERM algorithm for machine learning is proposed by Li et al. (2021), and good performance of the TERM under outlier and noisy label scenarios for negative tilting ($\gamma<0$) and under imbalance and fairness constraints for positive tilting ($\gamma>0$) is demonstrated. Inspired by TERM, Wang et al. (2023) propose a class of new tilted sparse additive models based on the Tikhonov regularization scheme. Their results have some limitations. First, in (Wang et al., 2023, Theorem 3.3) the authors derive an upper bound for $\lambda=n^{-\zeta}$ where $\zeta<-1/2$ and $\lambda$ are the regularization parameters in (Wang et al., 2023, Eq.4). This implies $\lambda\rightarrow\infty$ as $n\rightarrow+\infty$, which is impractical. Finally, the analysis in (Wang et al., 2023) assumes that both the loss function and its derivative are bounded. Therefore, it can not be applied to the unbounded loss function scenario. Furthermore, we consider KL regularization, which is different from the Tikhonov regularization scheme with the sparsity-induced $\ell_{1,2}$-norm regularizer as introduced in (Wang et al., 2023). Therefore, our current results do not cover the learning algorithm in [9]. Zhang et al. (2023) studied the TERM as a target function to improve the robustness of estimators. The application of TERM in federated learning is also studied, in (Li et al., 2023b; Zhang et al., 2022). The authors in (Lee et al., 2020), propose an upper bound on the Entropic risk function generalization error via the representation of coherent risk function and using the Rademacher complexity approach. However, their approach is limited to negative tilt and bounded loss function. Although rich experiments are given by Li et al. (2021) for the TERM algorithm in different applications, the generalization error of the TERM has not yet been addressed for unbounded loss functions.
Unbounded loss functions: Some works studied the generalization error under unbounded loss functions via the PAC-Bayesian approach. Losses with heavier tails are studied by Alquier and Guedj (2018) where probability bounds (non-high probability) are developed. Using a different estimator than empirical risk, PAC-Bayes bounds for losses with bounded second and third moments are developed by Holland (2019). Notably, their bounds include a term that can increase with the number of samples $n$. Kuzborskij and Szepesvári (2019) and Haddouche and Guedj (2022) also provide bounds for losses with a bounded second moment. The bounds in Haddouche and Guedj (2022) rely on a parameter that must be selected before the training data is drawn. Information-theoretic bounds based obased on the second moment of $\sup_{h\in\mathcal{H}}|\ell(h,Z)-\mathbb{E}[\ell(h,\tilde{Z})]|$ are derived in Lugosi and Neu (2022, 2023). In contrast, our second moment assumption is more relaxed, being based on the expected version with respect to the distribution over the hypothesis set and the data-generating distribution. In our paper, we provided a generalization error bound on the tilted empirical risk via uniform and information-theoretic approaches under a bounded second-moment assumption.
Appendix C Technical Tools
We first define the functional linear derivative as in Cardaliaguet et al. (2019).
Definition C.1.
(Cardaliaguet et al., 2019) A functional $U:\mathcal{P}(\mathbb{R}^{n})\to\mathbb{R}$ admits a functional linear derivative if there is a map $\frac{\delta U}{\delta m}:\mathcal{P}(\mathbb{R}^{n})\times\mathbb{R}^{n}\to\mathbb{R}$ which is continuous on $\mathcal{P}(\mathbb{R}^{n})$, such that for all $m,m^{\prime}\in\mathcal{P}(\mathbb{R}^{n})$, it holds that
$U(m^{\prime})-U(m)=\!\int_{0}^{1}\int_{\mathbb{R}^{n}}\frac{\delta U}{\delta m}(m_{\lambda},a)\,(m^{\prime}-m)(da)\,\mathrm{d}\lambda,$
where $m_{\lambda}=m+\lambda(m^{\prime}-m)$.
The following lemmas are used in our proofs.
Lemma C.2.
Suppose that $X>0$ and $\gamma<0$, then we have
$\mathrm{Var}(\exp(\gamma X))\leq\gamma^{2}\mathrm{Var}(X).$ (32)
Proof.
By the mean value theorem, for each realisation $X(\omega)$ of $X$ for an element $\omega$ of its underlying probability space there is a value $c(\omega)$ in the interval between $X(\omega)$ and $\mathbb{E}[X]$ such that
$\exp(\gamma X(\omega))-\exp(\gamma\mathbb{E}[X])=\gamma(X-\mathbb{E}[X])\exp(\gamma c(\omega)).$
As $X>0$ we have $c(\omega)>0$. Moreover,
$\begin{split}\mathrm{Var}(\exp(\gamma X))&=\mathbb{E}[(\exp(\gamma X)-\mathbb{E}[\exp(\gamma X)])^{2}]\\ &\overset{(a)}{\leq}\mathbb{E}[(\exp(\gamma X)-\exp(\gamma\mathbb{E}[X]))^{2}]\\ &\overset{(b)}{=}\mathbb{E}[\gamma^{2}\exp(2\gamma c)(X-\mathbb{E}[X])^{2}]\\ &\overset{(c)}{\leq}\gamma^{2}\mathrm{Var}(X),\end{split}$
where $(a)$, $(b)$ and $(c)$ follow from the minimum mean square representation, the mean-value theorem and the negativity of $\gamma X$, respectively. ∎
Lemma C.3.
Suppose that $0<a<X<b<\infty$. Then the following inequality holds,
$\begin{split}\frac{\mathrm{Var}_{P_{X}}(X)}{2b^{2}}\leq\log(\mathbb{E}[X])-\mathbb{E}[\log(X)]\leq\frac{\mathrm{Var}_{P_{X}}(X)}{2a^{2}},\end{split}$
where $\mathrm{Var}_{P_{X}}(X)$ is the variance of $X$ under the distribution $P_{X}$.
Proof.
As $\frac{\mathrm{d}^{2}}{\mathrm{d}x^{2}}\left(\log(x)+\beta x^{2}\right)=\frac{-1}{x^{2}}+2\beta,$ the function $\log(x)+\beta x^{2}$ is concave for $\beta=\frac{1}{2b^{2}}$ and convex for $\beta=\frac{1}{2a^{2}}$. Hence, by Jensen’s inequality,
$\begin{split}\mathbb{E}[\log(X)]&=\mathbb{E}\Big{[}\log(X)+\frac{X^{2}}{2b^{2}}-\frac{X^{2}}{2b^{2}}\Big{]}\\ &\leq\log(\mathbb{E}[X])+\frac{1}{2b^{2}}\mathbb{E}[X]^{2}-\frac{1}{2b^{2}}\mathbb{E}[X^{2}]\\ &=\log(\mathbb{E}[X])-\frac{1}{2b^{2}}\mathrm{Var}_{P_{X}}(X),\end{split}$
which completes the proof of the lower bound. A similar approach can be applied to derive the upper bound. ∎
Lemma C.4.
Suppose that $0<a<X<b<\infty$. Then the following inequality holds,
$\begin{split}\frac{-\mathrm{Var}_{P_{X}}(X)\exp(b)}{2}\leq\exp(E[X])-E[\exp(X)]\leq\frac{-\mathrm{Var}_{P_{X}}(X)\exp(a)}{2},\end{split}$
where $\mathrm{Var}_{P_{X}}(X)$ is the variance of $X$ under the distribution $P_{X}$.
In the next results, $P_{S}$ is the distribution of $S$.
Lemma C.5 (McDiarmid’s inequality (McDiarmid, 1989)).
Let $F:\mathcal{Z}^{n}\mapsto\mathbb{R}$ be any measurable function for which there exists constants $c_{i},i=1,\cdots,m)$ such that
$\sup_{S\in\mathcal{Z}^{n},\tilde{Z}_{i}\in\mathcal{Z}}|F(S)-F(S_{(i)}|\leq c_{i},$
where $S_{(i)}=\{Z_{1},\cdots,\tilde{Z}_{i},\cdots,Z_{n}\},$ is the replace-one sample dataset and $\tilde{Z}_{i}$ is an i.i.d. sample with respect to all $Z_{i}$ for $i\in[n]$. Then the following inequality holds with probability at least $(1-\delta)$ under $P_{S}$,
$|F(S)-\mathbb{E}_{P_{S}}[F(S)]|\leq\sqrt{\big{(}\sum_{i=1}^{n}c_{i}^{2}\big{)}\frac{\log(1/\delta)}{2}}.$
Lemma C.6 (Hoeffding Inequality, Boucheron et al., 2013).
Suppose that $S=\{Z_{i}\}_{i=1}^{n}$ are bounded independent random variables such that $a\leq Z_{i}\leq b,i=1,\ldots,n$. Then the following inequality holds with probability at least $(1-\delta)$ under $P_{S}$,
$\Big{|}\mathbb{E}[Z]-\frac{1}{n}\sum_{i=1}^{n}Z_{i}\Big{|}\leq(b-a)\sqrt{\frac{\log(2/\delta)}{2n}}.$ (33)
Lemma C.7 (Bernstein’s Inequality, Boucheron et al., 2013).
Suppose that $S=\{Z_{i}\}_{i=1}^{n}$ are i.i.d. random variable such that $|Z_{i}-\mathbb{E}[Z]|\leq R$ almost surely for all $i$, and $\mathrm{Var}(Z)=\sigma^{2}$. Then the following inequality holds with probability at least $(1-\delta)$ under $P_{S}$,
$\Big{|}\mathbb{E}[Z]-\frac{1}{n}\sum_{i=1}^{n}Z_{i}\Big{|}\leq\sqrt{\frac{4\sigma^{2}\log(2/\delta)}{n}}+\frac{4R\log(2/\delta)}{3n}.$ (34)
Lemma C.8.
For a positive random variable, $Z>0$, suppose $\mathbb{E}[Z^{2}]\leq\eta$ Then, the following inequality holds,
$\mathbb{E}[Z]\leq\eta^{1/2}.$
Lemma C.9.
Suppose $\mathbb{E}[X^{2}]<\infty$. Then, for $X>0$ and $\gamma<0$ the following inequality holds,
$0\leq\mathbb{E}[X]-\frac{1}{\gamma}\log{\mathbb{E}[e^{\gamma X}]}\leq\frac{-\gamma}{2}\mathbb{E}[X^{2}].$
Proof.
The left inequality follows from Jensen’s inequality applied to $f(x)=\log{(x)}$. For the right inequality, we have for $\gamma X<0$,
$e^{\gamma X}\leq 1+\gamma X+\frac{1}{2}(\gamma X)^{2}.$ (35)
Therefore, we have
$\begin{split}\frac{1}{\gamma}\log{\mathbb{E}[e^{\gamma X}]}&\geq\frac{1}{\gamma}\log{\mathbb{E}\left[1+\gamma X+\frac{1}{2}\gamma^{2}X^{2}\right]}\\ &=\frac{1}{\gamma}\log{\left(1+\gamma\mathbb{E}[X]+\frac{1}{2}\gamma^{2}\mathbb{E}[X^{2}]\right)}\\ &\geq\frac{1}{\gamma}\left(\gamma\mathbb{E}[X]+\frac{1}{2}\gamma^{2}\mathbb{E}[X^{2}]\right)\\ &=\mathbb{E}[X]+\frac{\gamma}{2}\mathbb{E}[X^{2}].\end{split}$
Lemma C.10.
Suppose that $0<a<x<b$ and $f(x)$ is an increasing and concave function. Then the following holds,
$f^{\prime}(b)(b-a)\leq f(b)-f(a)\leq f^{\prime}(a)(b-a).$ (36)
Lemma C.11 (Uniform bound (Mohri et al., 2018)).
Let $\mathcal{F}$ be the set of functions $f:\mathcal{Z}\to[0,M]$ and $\mu$ be a distribution over $\mathcal{Z}$. Let $S=\{z_{i}\}_{i=1}^{n}$ be a set of size $n$ i.i.d. drawn from $\mathcal{Z}$. Then, for any $\delta\in(0,1)$, with probability at least $1-\delta$ over the choice of $S$, we have
$\sup_{f\in\mathcal{F}}\left\{\mathbb{E}_{Z\sim\mu}[f(Z)]-\frac{1}{n}\sum_{i=1}^{n}f(z_{i})\right\}\leq 2\hat{\mathfrak{R}}_{S}(\mathcal{F})+3M\sqrt{\frac{1}{2n}\log\frac{2}{\delta}}.$
We use the next two results, namely Talagrand’s contraction lemma and Massart’s Lemma, to estimate the Rademacher complexity.
Lemma C.12 (Talagrand’s contraction lemma (Shalev-Shwartz and Ben-David, 2014)).
Let $\phi_{i}:\mathbb{R}\rightarrow\mathbb{R}$ $(i\in\{1,\ldots,n\})$ be $L$-Lipschitz functions and $\mathcal{F}_{r}$ be a set of functions from $\mathcal{Z}$ to $\mathbb{R}$. Then it follows that for any $\{z_{i}\}_{i=1}^{n}\subset\mathcal{Z}$,
$\mathbb{E}_{\sigma}\left[\sup_{f\in\mathcal{F}_{r}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi_{i}(f(z_{i}))\right]\leq L\mathbb{E}_{\sigma}\left[\sup_{f\in\mathcal{F}_{r}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}f(z_{i})\right].$
Lemma C.13 (Massart’s lemma (Massart, 2000)).
Assume that the hypothesis space $\mathcal{H}$ is finite. Let $B^{2}:=\max_{h\in\mathcal{H}}\Big{(}\sum_{i=1}^{n}h^{2}(z_{i})\Big{)}.$ Then
^$\hat{\mathfrak{R}}_{S}(\mathcal{H})\leq\frac{B\sqrt{2\log(\mathrm{card}(\mathcal{H}))}}{n}.$
Lemma C.14 (Lemma 1 in (Xu and Raginsky, 2017)).
For any measureable function $f:\mathcal{Z}\rightarrow[0,M]$,
$\Big{|}\mathbb{E}_{P_{X,Y}}[f(X,Y)]-\mathbb{E}_{P_{X}\otimes P_{Y}}[f(X,Y)]\Big{|}\leq M\sqrt{\frac{I(X;Y)}{2}}.$
Lemma C.15 (Coupling Lemma).
Assume the function $f:\mathcal{Z}\rightarrow[0,M]$ and the function $g:\mathbb{R}^{+}\mapsto\mathbb{R}^{+}$ are $L$-Lipschitz. Then the following upper bound holds,
$\Big{|}\mathbb{E}_{P_{X,Y}}[g\circ f(X,Y)]-\mathbb{E}_{P_{X}\otimes P_{Y}}[g\circ f(X,Y)]\Big{|}\leq LM\sqrt{\frac{I(X;Y)}{2}}.$
We recall that the notation $\mathbb{TV}$ denotes the total variation distance between probability distributions.
Lemma C.16 (Kantorovich-Rubenstein duality of total variation distance, see  (Polyanskiy and Wu, 2022)).
The Kantorovich-Rubenstein duality (variational representation) of the total variation distance is as follows:
$\mathbb{TV}(m_{1},m_{2})=\frac{1}{2L}\sup_{g\in\mathcal{G}_{L}}\left\{\mathbb{E}_{Z\sim m_{1}}[g(Z)]-\mathbb{E}_{Z\sim m_{2}}[g(Z)]\right\},$ (37)
where $\mathcal{G}_{L}=\{g:\mathcal{Z}\rightarrow\mathbb{R},||g||_{\infty}\leq L\}$.
Appendix D Proofs and Details of Section 3
D.1 Uniform bound: details for bounded loss
Proposition 3.2 (Restated).
Under Assumption 3.1, the difference between the population risk and the tilted population risk for $\gamma\in\mathbb{R}$, satisfies that for any $h\in\mathcal{H}$
$\begin{split}\frac{-1}{2\gamma}\mathrm{Var}(\exp(\gamma\ell(h,Z)))\leq\mathrm{R}(h,\mu)-\mathrm{R}_{\gamma}(h,\mu^{\otimes n})\leq\frac{-\exp(-2\gamma M)}{2\gamma}\mathrm{Var}(\exp(\gamma\ell(h,Z))).\end{split}$
Proof.
For any $h\in\mathcal{H}$ we have
$\begin{split}\mathrm{R}(h,\mu)&=\mathbb{E}[\ell(h,Z)]\\ &=\mathbb{E}\Big{[}\frac{1}{\gamma}\log\big{(}\exp(\gamma\ell(h,Z))\big{)}\Big{]}\\ &=\frac{1}{|\gamma|}\Bigg{[}\mathbb{E}_{Z\sim\mu}\Big{[}-\log\big{(}\exp(\gamma\ell(h,Z))\big{)}-\frac{\exp(-2\gamma M)}{2}\exp(2\gamma\ell(h,Z))\\ &\qquad+\frac{\exp(-2\gamma M)}{2}\exp(2\gamma\ell(h,Z))\Big{]}\Bigg{]}\\ &\leq\frac{1}{\gamma}\log(\mathbb{E}[\exp(\gamma\ell(h,\mu))])+\frac{\exp(-2\gamma M)}{2|\gamma|}\mathrm{Var}(\exp(\gamma\ell(h,Z)))\\ &=\mathrm{R}_{\gamma}(h,\mu^{\otimes n})+\frac{\exp(-2\gamma M)}{2|\gamma|}\mathrm{Var}(\exp(\gamma\ell(h,Z))).\end{split}$ (38)
A similar approach can be applied for $\gamma>0$ by using Lemma C.3 and the final result holds. ∎
Theorem 3.3.
Given Assumption 3.1, for any fixed $h\in\mathcal{H}$ with probability at least $(1-\delta)$ the tilted generalization error satisfies the upper bound, $\begin{split}\mathrm{gen}_{\gamma}(h,S)\leq\frac{-\exp(-2\gamma M)}{2\gamma}\operatorname{Var}\big{(}\exp(\gamma\ell(h,Z))\big{)}+\frac{\big{(}\exp(|\gamma|M)-1\big{)}}{|\gamma|}\sqrt{\frac{\log(2/\delta)}{2n}}.\end{split}$ (39)
Proof.
We can apply the Proposition 3.2 to provide an upper bound on term $I_{1}$. Regarding the term $I_{5}$, we have for $\gamma<0$
$\begin{split}&\mathrm{R}_{\gamma}(h,\mu^{\otimes n})-\widehat{\mathrm{R}}_{\gamma}(h,S)\\ &=\frac{1}{\gamma}\log(\mathbb{E}_{\mu^{\otimes n}}[\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(h,Z_{i}))])-\frac{1}{\gamma}\log(\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(h,Z_{i})))\\ &\leq\frac{\exp(-\gamma M)}{|\gamma|}\big{|}\mathbb{E}_{\tilde{Z}\sim\mu}[\exp(\gamma\ell(h,\tilde{Z}))]-\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(h,Z_{i}))\big{|}\\ &\leq\frac{\exp(-\gamma M)(1-\exp(\gamma M))}{|\gamma|}\sqrt{\frac{\log(2/\delta)}{2n}}.\end{split}$ (40)
Similarly, for $\gamma>0$, we have
$\mathrm{R}_{\gamma}(h,\mu^{\otimes n})-\widehat{\mathrm{R}}_{\gamma}(h,S)\leq\frac{(\exp(\gamma M)-1)}{|\gamma|}\sqrt{\frac{\log(2/\delta)}{2n}}.$ (41)
Combining this bound with Proposition 3.2 completes the proof. ∎
Corollary 3.7 (Restated).
Under the same Assumptions as in Theorem 3.3 and assuming $\gamma$ is of order $O(n^{-\beta})$ for $\beta>0$, the upper bound on the tilted generalization error in Theorem 3.3 has a convergence rate of $\max\big{(}O(1/\sqrt{n}),O(n^{-\beta})\big{)}$ as $n\rightarrow\infty$.
Proof.
Using the inequality $\frac{x}{x+1}\leq\log(1+x)\leq x$ and Taylor expansion for the exponential function,
$\exp(|\gamma|M)=1+|\gamma|M+\frac{|\gamma|^{2}M^{2}}{2}+\frac{|\gamma|^{3}M^{3}}{6}+O(\gamma^{4}),$ (42)
it follows that
$\begin{split}&\frac{\big{(}\exp(|\gamma|M)-1\big{)}}{|\gamma|}\approx M+\frac{|\gamma|M^{2}}{2}+\frac{|\gamma|^{2}M^{3}}{6}+O(\gamma^{3}).\end{split}$ (43)
This results in a convergence rate of $O(1/\sqrt{n})$ for $\frac{\big{(}\exp(|\gamma|M)-1\big{)}}{|\gamma|}\sqrt{\frac{\log(\mathrm{card}(\mathcal{H}))+\log(2/\delta)}{2n}}$ under $\gamma\rightarrow 0$.
For the term $\frac{\max(1,\exp(-2\gamma M))(1-\exp(\gamma M))^{2}}{8|\gamma|}$, using Taylor expansion, we have the convergence rate of $O(|\gamma|)$ for the first term; this completes the proof. ∎
Theorem 3.4.
Under the same assumptions of Theorem 3.3, for a fixed $h\in\mathcal{H}$, with probability at least $(1-\delta)$, the tilted generalization error satisfies the lower bound $\begin{split}\mathrm{gen}_{\gamma}(h,S)\geq\frac{-1}{2\gamma}\operatorname{Var}\big{(}\exp(\gamma\ell(h,Z))\big{)}-\frac{\big{(}\exp(|\gamma|M)-1\big{)}}{|\gamma|}\sqrt{\frac{\log(2/\delta)}{2n}}.\end{split}$ (44)
Proof.
The proof is similar to that of Theorem 3.3, by using the lower bound in Proposition 3.2. ∎
Corollary 3.6.
Let $A(\gamma)=(1-\exp(\gamma M))^{2}$. Under the same assumptions in Theorem 3.3, with probability at least $(1-\delta)$, and a finite hypothesis space, the absolute value of the titled generalization error satisfies $\begin{split}\sup_{h\in\mathcal{H}}|\mathrm{gen}_{\gamma}(h,S)|&\leq\frac{\big{(}\exp(|\gamma|M)-1\big{)}}{|\gamma|}\sqrt{\frac{\log(\mathrm{card}(\mathcal{H}))+\log(2/\delta)}{2n}}+\frac{\max(1,\exp(-2\gamma M))A(\gamma)}{8|\gamma|},\end{split}$ where $A(\gamma)=(1-\exp(\gamma M))^{2}$.
Proof.
We can derive the following upper bound on the absolute of tilted generalization error by combining Theorem 3.3 and Theorem 3.4 for any fixed $h\in\mathcal{H}$
$\begin{split}|\mathrm{gen}_{\gamma}(h,S)|&\leq\frac{\big{(}\exp(|\gamma|M)-1\big{)}}{|\gamma|}\sqrt{\frac{\log(\mathrm{card}(\mathcal{H}))+\log(2/\delta)}{2n}}+\frac{\max(1,\exp(-2\gamma M))A(\gamma)}{8|\gamma|},\end{split}$ (45)
where $A(\gamma)=(1-\exp(\gamma M))^{2}$. Then, the final result follow by applying the uniform bound for all $h\in\mathcal{H}$ using (45). ∎
Corollary 3.11 (Restated).
Under the same assumptions in Theorem 3.3, and a finite hypothesis space, with probability at least $(1-\delta)$, the excess risk of tilted empirical risk satisfies $\begin{split}\mathfrak{E}_{\gamma}(\mu)&\leq\frac{2\big{(}\exp(|\gamma|M)-1\big{)}}{|\gamma|}\sqrt{\frac{\log(\mathrm{card}(\mathcal{H}))+\log(2/\delta)}{2n}}+\frac{2\max(1,\exp(-2\gamma M))A(\gamma)}{8|\gamma|},\end{split}$ where $A(\gamma)=(1-\exp(\gamma M))^{2}$.
Proof.
It can be proved that,
$\mathfrak{E}_{\gamma}(\mu)\leq 2\sup_{h\in\mathcal{H}}|\mathrm{gen}_{\gamma}(h,S)|$
$\mathrm{R}(h_{\gamma}^{*}(S),\mu)$ $\leq\hat{\mathrm{R}}_{\gamma}(h_{\gamma}^{*}(S),\mu)+U$ $\leq\hat{\mathrm{R}}_{\gamma}(h^{*}(\mu),\mu)+U$ $\leq\mathrm{R}(h^{*}(\mu),\mu)+2U,$
where $U=\sup_{h\in\mathcal{H}}|\mathrm{R}(h,\mu)-\hat{\mathrm{R}}_{\gamma}(h,S)|=\sup_{h\in\mathcal{H}}|\operatorname{gen}_{\gamma}(h,S)|.$
Note that $\sup_{h\in\mathcal{H}}|\operatorname{gen}_{\gamma}(h,S)|$ can be bounded using Corollary 3.6. ∎
D.1.1 Another Approach for Uniform Bounds
For this approach, we decompose the tilted generalization error as follows,
$\begin{split}&\mathrm{gen}_{\gamma}(h,S)\\ &=\underbrace{\mathrm{R}(h,\mu)-\mathrm{R}_{\gamma}(h,\mu^{\otimes n})}_{I_{1}}+\underbrace{\mathrm{R}_{\gamma}(h,\mu^{\otimes n})-\overline{\mathrm{R}}_{\gamma}(h,\mu^{\otimes n})}_{I_{2}}+\underbrace{\overline{\mathrm{R}}_{\gamma}(h,\mu^{\otimes n})-\widehat{\mathrm{R}}_{\gamma}(h,S)}_{I_{3}},\end{split}$
where $I_{1}$ the same as in (12), $I_{2}$ is the difference between the tilted population risk and the expected TER, and $I_{3}$ is the difference between the expected TER and the TER.
Proposition D.1.
Under Assumption 3.1, for all $h\in\mathcal{H}$, for $\gamma\in\mathbb{R}\setminus\{0\}$, we have
$\frac{\exp(-2\gamma M)}{2\gamma n}\mathrm{Var}\big{(}\exp(\gamma\ell(h,Z))\big{)}\leq\mathrm{R}_{\gamma}(h,\mu^{\otimes n})-\overline{\mathrm{R}}_{\gamma}(h,\mu^{\otimes n})\leq\frac{1}{2\gamma n}\mathrm{Var}\big{(}\exp(\gamma\ell(h,Z))\big{)}.$
Proof.
Due to the i.i.d. assumption, we have
$\begin{split}\mathrm{R}_{\gamma}(h,\mu^{\otimes n})&=\frac{1}{\gamma}\log(\mathbb{E}_{Z\sim\mu}[\exp(\gamma\ell(h,Z))])\\ &=\frac{1}{\gamma}\log\Bigg{(}\mathbb{E}_{S\sim\mu^{\otimes n}}\bigg{[}\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(h,Z_{i}))\bigg{]}\Bigg{)}\end{split}$
Using Lemma C.3 and Jensen’sinequality, we have for $\gamma<0$,
$\begin{split}&\overline{\mathrm{R}}_{\gamma}(h,\mu^{\otimes n})\\ &=\frac{1}{\gamma}\mathbb{E}\Bigg{[}\log\Bigg{(}\bigg{[}\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(h,Z_{i}))\bigg{]}\Bigg{)}\Bigg{]}\\ &=\frac{1}{|\gamma|}\mathbb{E}_{S\sim\mu^{\otimes n}}\Bigg{[}-\log\Bigg{(}\bigg{[}\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(h,Z_{i}))\bigg{]}\Bigg{)}-\frac{\exp(-2\gamma M)}{2}\bigg{[}\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(h,Z_{i}))\bigg{]}^{2}\\ &\qquad+\frac{\exp(-2\gamma M)}{2}\bigg{[}\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(h,Z_{i}))\bigg{]}^{2}\Bigg{]}\\ &\leq\frac{1}{\gamma}\log\big{(}\mathbb{E}_{Z\sim\mu}[\exp(\gamma\ell(h,Z))]\big{)}+\frac{\exp(-2\gamma M)}{2|\gamma|}\mathrm{Var}\Big{(}\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(h,Z_{i}))\Big{)}\\ &=\mathrm{R}_{\gamma}(h,\mu^{\otimes n})+\frac{\exp(-2\gamma M)}{2|\gamma|n}\mathrm{Var}_{Z\sim\mu}\Big{(}\exp(\gamma\ell(h,Z))\Big{)}.\end{split}$ (46)
Therefore, we have
$\overline{\mathrm{R}}_{\gamma}(h,\mu^{\otimes n})-\mathrm{R}_{\gamma}(h,\mu^{\otimes n})\leq\frac{\exp(-2\gamma M)}{2|\gamma|n}\mathrm{Var}_{Z\sim\mu}\Big{(}\exp(\gamma\ell(h,Z))\Big{)}.$
Similarly, we can show that,
$\overline{\mathrm{R}}_{\gamma}(h,\mu^{\otimes n})-\mathrm{R}_{\gamma}(h,\mu^{\otimes n})\geq\frac{1}{2|\gamma|n}\mathrm{Var}_{Z\sim\mu}\Big{(}\exp(\gamma\ell(h,Z))\Big{)}.$
Thus, for $\gamma<0$ we obtain
$\frac{1}{2\gamma n}\mathrm{Var}_{Z\sim\mu}\Big{(}\exp(\gamma\ell(h,Z))\Big{)}$$\geq$ $\mathrm{R}_{\gamma}(h,\mu^{\otimes n})-\overline{\mathrm{R}}_{\gamma}(h,\mu^{\otimes n})$
$\geq$ $\frac{\exp(-2\gamma M)}{2\gamma n}\mathrm{Var}_{Z\sim\mu}\Big{(}\exp(\gamma\ell(h,Z))\Big{)}.$
The same approach can be applied to $\gamma>0$. This completes the proof. ∎
Next we use McDiarmid’s inequality, Lemma C.5, to derive an upper bound on the absolute value of the term $I_{3}$, in the following proposition.
Proposition D.2.
Under Assumption 3.1 and assuming $|\gamma|<\frac{\log(n+1)}{M}$, the difference of expected TER and TER satisfies with probability at least $(1-\delta)$,
$\begin{split}|\overline{\mathrm{R}}_{\gamma}(h,\mu^{\otimes n})-\widehat{\mathrm{R}}_{\gamma}(h,S)|\leq c\sqrt{\frac{n\log(1/\delta)}{2}},\end{split}$ (47)
where $c=\Big{|}\frac{1}{\gamma}\log\big{(}1+\frac{1-\exp(|\gamma|M)}{n}\big{)}\Big{|}.$
Proof.
Viewing TER as a function of the data samples $\{Z_{i}\}_{i=1}^{n}$, the following bounded difference holds for $\gamma<0$, with the supremum taken over $\{z_{i}\}_{i=1}^{n}\in\mathcal{Z}^{n},\tilde{z}_{i}\in\mathcal{Z}$,
$\begin{split}&\sup|\widehat{\mathrm{R}}_{\gamma}(h,S)-\widehat{\mathrm{R}}_{\gamma}(h,S_{(i)})|\\ &=\sup\Big{|}\frac{1}{\gamma}\Big{(}\log(\frac{1}{n}\sum_{j=1}^{n}\exp(\gamma\ell(h,z_{j})))-\log(\frac{1}{n}\sum_{\begin{subarray}{c}j=1,\\ j\neq i\end{subarray}}^{n}\exp(\gamma\ell(h,z_{j}))+\exp(\gamma\ell(h,\tilde{z}_{i})))\Big{)}\Big{|}\\ &=\sup\Big{|}\frac{1}{\gamma}\Big{(}\log\Bigg{(}\frac{\sum_{j=1}^{n}\exp(\gamma\ell(h,z_{j})))}{\sum_{\begin{subarray}{c}j=1,\\ j\neq i\end{subarray}}^{n}\exp(\gamma\ell(h,z_{j}))+\exp(\gamma\ell(h,\tilde{z}_{i}))}\Bigg{)}\Big{)}\Big{|}\\ &=\sup\Big{|}\frac{1}{\gamma}\log\Bigg{(}1+\frac{\exp(\gamma\ell(h,\tilde{z}_{i}))-\exp(\gamma\ell(h,z_{i}))}{\sum_{\begin{subarray}{c}j=1\end{subarray}}^{n}\exp(\gamma\ell(h,z_{j}))}\Bigg{)}\Big{|}\\ &\leq\max\Bigg{(}\Big{|}\frac{1}{\gamma}\log\Bigg{(}1+\frac{\exp(\gamma M)-1}{n\exp(\gamma M)}\Bigg{)}\Big{|},\Big{|}\frac{1}{\gamma}\log\Bigg{(}1+\frac{1-\exp(\gamma M)}{n\exp(\gamma M)}\Bigg{)}\Big{|}\Bigg{)}\\ &=\max\Bigg{(}\Big{|}\frac{1}{\gamma}\log\Bigg{(}1+\frac{\exp(-\gamma M)-1}{n}\Bigg{)}\Big{|},\Big{|}\frac{1}{\gamma}\log\Bigg{(}1+\frac{1-\exp(-\gamma M)}{n}\Bigg{)}\Big{|}\Bigg{)}.\end{split}$ (48)
Therefore, we choose
$\tilde{c}_{i}=\max\Bigg{(}\Big{|}\frac{1}{\gamma}\log\Bigg{(}1+\frac{\exp(-\gamma M)-1}{n}\Bigg{)}\Big{|},\Big{|}\frac{1}{\gamma}\log\Bigg{(}1+\frac{1-\exp(-\gamma M)}{n}\Bigg{)}\Big{|}\Bigg{)}$
in McDiarmid’s inequality, Lemma C.5. Note that for the term $\frac{1}{\gamma}\log\Bigg{(}1+\frac{1-\exp(-\gamma M)}{n}\Bigg{)}$, we need that $\frac{1-\exp(-\gamma M)}{n}>-1$, and therefore we should have $|\gamma|<\frac{\log(n+1)}{M}$; otherwise $\tilde{c}=\infty$.
Similarly, for $\gamma>0$, we can show that,
$\tilde{c}_{i}=\max\Bigg{(}\Big{|}\frac{1}{\gamma}\log\Bigg{(}1+\frac{\exp(\gamma M)-1}{n}\Bigg{)}\Big{|},\Big{|}\frac{1}{\gamma}\log\Bigg{(}1+\frac{1-\exp(\gamma M)}{n}\Bigg{)}\Big{|}\Bigg{)}.$
Then, under the distribution $P_{S}$, with probability at least $(1-\delta)$, the following inequality holds,
$|\overline{\mathrm{R}}_{\gamma}(h,\mu^{\otimes n})-\widehat{\mathrm{R}}_{\gamma}(h,S)|\leq\tilde{c}\sqrt{\frac{n\log(1/\delta)}{2}},$ (49)
where $\tilde{c}=\max\Big{(}\Big{|}\frac{1}{\gamma}\log\big{(}1-y\big{)}\Big{|},\Big{|}\frac{1}{\gamma}\log\big{(}1+y\big{)}\Big{|}\Big{)},$ and $y=\frac{1-\exp(|\gamma|M)}{n}$.
As we have $|\log(1-x)|\leq|\log(1+x)|$ for $-1<x<0$, it holds that $\tilde{c}=\Big{|}\frac{1}{\gamma}\log\big{(}1+y\big{)}\Big{|},$ and $y=\frac{1-\exp(|\gamma|M)}{n}$. ∎
With Propositions  D.23.2, we derive upper and lower bounds on the tilted generalization error.
Theorem D.3.
Given Assumption 3.1, $|\gamma|<\frac{\log(n+1)}{M}$, and assuming a finite hypothesis space, then for $\gamma\in\mathbb{R}\setminus\{0\}$ we have for a fixed $h\in\mathcal{H}$
$\begin{split}\mathrm{gen}_{\gamma}(h,S)\leq\Big{|}\frac{(1-\exp(\gamma M))^{2}}{8\gamma}\Big{(}\frac{1}{n}-\exp(-2\gamma M)\Big{)}\Big{|}+c\sqrt{\frac{n(\log(1/\delta))}{2}},\end{split}$ (50)
where $c=c(\gamma)=\Big{|}\frac{1}{\gamma}\log\big{(}1+\frac{1-\exp(|\gamma|M)}{n}\big{)}\Big{|}.$
Proof.
Combining the results in Proposition D.1, Proposition D.2 and Proposition 3.2, we can derive the following upper bound on generalization error,
$\begin{split}\mathrm{gen}_{\gamma}(h,S)\leq&\mathrm{Var}_{Z\sim\mu}(\exp(\gamma\ell(h,Z)))\Big{(}\frac{1}{2\gamma n}-\frac{\exp(-2\gamma M)}{2\gamma}\Big{)}\\ &+c\sqrt{\frac{n(\log(\mathrm{card}(\mathcal{H}))+\log(1/\delta))}{2}}.\end{split}$ (51)
From the boundedness of $\exp(\gamma\ell(h,z))\in[\exp(\gamma M),1]$ for $\gamma<0$ and $\exp(\gamma\ell(h,z))\in[1,\exp(\gamma M)]$ for $\gamma>0$, we have
$\mathrm{Var}_{Z\sim\mu}\Big{(}\exp(\gamma\ell(h,Z))\Big{)}\leq\frac{(1-\exp(\gamma M))^{2}}{4}.$
Substituting this in (51) completes the proof. ∎
Remark D.4 (The influence of $\gamma$).
As $\gamma\rightarrow 0$, the upper bound in Theorem D.3 on the tilted generalization error converges to the upper bound on the generalization error under the ERM algorithm obtained by Shalev-Shwartz and Ben-David (2014), (15). In particular, $c=c(\gamma)\rightarrow\frac{M}{n}$ and the first term in (50) vanishes. Therefore, the upper bound converges to a uniform bound on the linear empirical risk.
Corollary D.5.
Under the same assumptions in Corollary D.5, the upper bound on the tilted generalization error in Theorem D.3 has a convergence rate of $\max\big{(}O(1/\sqrt{n}),O(n^{-\beta})\big{)}$ as $n\rightarrow\infty$.
Proof.
Using the inequality $\frac{x}{x+1}\leq\log(1+x)\leq x$ and Taylor expansion for the exponential function,
$\exp(|\gamma|M)=1+|\gamma|M+\frac{|\gamma|^{2}M^{2}}{2}+\frac{|\gamma|^{3}M^{3}}{6}+O(\gamma^{4}),$ (52)
it follows that
$\begin{split}|\overline{\mathrm{R}}_{\gamma}(h,\mu^{\otimes n})-\widehat{\mathrm{R}}_{\gamma}(h,S)|&\leq\Big{|}\frac{\frac{1-\exp(|\gamma|M)}{n}}{|\gamma|(1+\frac{1-\exp(|\gamma|M))}{n}}\Big{|}\sqrt{\frac{n\log(1/\delta)}{2}}\\ &\leq\Big{|}\frac{1-\exp(|\gamma|M)}{|\gamma|(n+1-\exp(|\gamma|M))}\Big{|}\sqrt{\frac{n\log(1/\delta)}{2}}\\ &\leq\frac{M+(1/2)|\gamma|M^{2}+O(|\gamma|^{2})}{n-|\gamma|M-1/2\gamma^{2}M^{2}-O(|\gamma|^{3})}\sqrt{\frac{n\log(1/\delta)}{2}}.\end{split}$
This results in a convergence rate of $O(1/\sqrt{n})$ for $\gamma\rightarrow 0$.
For the term $\frac{(1-\exp(\gamma M))^{2}}{8\gamma}\Big{(}\frac{1}{n}-\exp(-2\gamma M)\Big{)}$, using the Taylor expansion (52), we have $O(\frac{\gamma}{n})$ for $\frac{(1-\exp(\gamma M))^{2}}{8n\gamma}$ and $O(\gamma)$ for $\frac{-(1-\exp(\gamma M))^{2}\exp(-2\gamma M)}{8\gamma}$. Therefore, the final result follows. ∎
Theorem D.6 (Uniform Lower Bound).
Under the assumptions of Theorem D.3 for a fixed $h\in\mathcal{H}$ we have
$\begin{split}\mathrm{gen}_{\gamma}(h,S)\geq\frac{\mathrm{Var}\big{(}\exp(\gamma\ell(h,Z))\big{)}}{2\gamma}\Big{(}\frac{\exp(-2\gamma M)}{n}-1\Big{)}-c\sqrt{\frac{n(\log(1/\delta))}{2}},\end{split}$ (53)
where $c=c(\gamma)=\Big{|}\frac{1}{\gamma}\log\big{(}1+\frac{1-\exp(|\gamma|M)}{n}\big{)}\Big{|}.$
Proof.
The final results ∎
Remark D.7.
For $\frac{\log(n)}{2M}<|\gamma|<\frac{\log(n+1)}{M}$, the lower bound in Theorem D.6 for a negative $\gamma$ can be tighter than the lower bound on the tilted generalization error, (53), for a positive $\gamma$, due to the term $\frac{\exp(-2\gamma M)}{n}-1$ in Theorem D.6.
Remark D.8.
The terms $\frac{(1/n-\exp(-2\gamma M))}{\gamma}$ in Theorem D.3 can cause the upper bounds in Theorem D.3 to be tighter for $\gamma>0$ than for $\gamma<0$. Furthermore, in comparison with the uniform upper bound on the linear generalization error, (15), we obtain a tilted generalization error upper bound, (50), which can be tighter, for sufficiently large value of samples $n$ and small values of $\gamma$.
D.2 Information-theoretic bounds: details for bounded loss
Proposition 3.12 (Restated).
Under Assumption 3.1, the following inequality holds,
$\begin{split}&\Big{|}\overline{\mathrm{R}}_{\gamma}(H,P_{H}\otimes\mu^{\otimes n})-\overline{\mathrm{R}}_{\gamma}(H,P_{H,S})\Big{|}\leq\frac{(\exp(|\gamma|M)-1)}{|\gamma|}\sqrt{\frac{I(H;S)}{2n}}.\end{split}$
Proof.
The proof follows directly from applying Lemma C.15 to the $\log(.)$ function and then applying Lemma C.14. ∎
Theorem 3.13 (Restated).
Under Assumption 3.1, the expected tilted generalization error satisfies $\overline{\mathrm{gen}}_{\gamma}(H,S)\leq\frac{(\exp(|\gamma|M)-1)}{|\gamma|}\sqrt{\frac{I(H;S)}{2n}}-\frac{\gamma\exp(-\gamma M)}{2}\big{(}1-\frac{1}{n}\big{)}\mathbb{E}_{P_{H}}\big{[}\mathrm{Var}_{\tilde{Z}\sim\mu}(\ell(H,\tilde{Z}))\big{]}.$
Proof.
We expand
$\begin{split}&\overline{\mathrm{gen}}_{\gamma}(H,S)=\mathrm{R}(H,P_{H}\otimes\mu)-\overline{\mathrm{R}}_{\gamma}(H,P_{H}\otimes\mu^{\otimes n})+\overline{\mathrm{R}}_{\gamma}(H,P_{H}\otimes\mu^{\otimes n})-\overline{\mathrm{R}}_{\gamma}(H,P_{H,S}).\end{split}$ (54)
Using Proposition 3.12, it follows that
$\begin{split}|\overline{\mathrm{R}}_{\gamma}(H,P_{H}\otimes\mu^{\otimes n})-\overline{\mathrm{R}}_{\gamma}(H,P_{H,S})|\leq\frac{(\exp(|\gamma|M)-1)}{|\gamma|}\sqrt{\frac{I(H;S)}{2n}}.\end{split}$ (55)
Using the Lipschitz property of the $\log(.)$ function under Assumption 3.1, we have for $\gamma>0$,
$\begin{split}&\mathrm{R}(H,P_{H}\otimes\mu)-\overline{\mathrm{R}}_{\gamma}(H,P_{H}\otimes\mu^{\otimes n})\\ &=\mathbb{E}_{P_{H}\otimes\mu^{\otimes n}}[\frac{1}{\gamma}\log(\exp(\frac{\gamma}{n}\sum_{i=1}^{n}\ell(H,Z_{i})))]-\mathbb{E}_{P_{H}\otimes\mu^{\otimes n}}[\frac{1}{\gamma}\log(\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i})))]\\ &\leq\frac{\exp(-\gamma M)}{\gamma}\mathbb{E}_{P_{H}\otimes\mu^{\otimes n}}\Big{[}\exp(\frac{\gamma}{n}\sum_{i=1}^{n}\ell(H,Z_{i}))-\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))\Big{]}\\ &\leq\frac{-\exp(-\gamma M)}{2\gamma}\mathbb{E}_{P_{H}\otimes\mu^{\otimes n}}\Bigg{[}\Big{(}\frac{1}{n}\sum_{i=1}^{n}\gamma^{2}\ell(H,Z_{i})^{2}\Big{)}-\Big{(}\frac{1}{n^{2}}\big{(}\sum_{i=1}^{n}\gamma\ell(H,Z_{i})\big{)}^{2}\Big{)}\Bigg{]}\\ &=\frac{-\exp(-\gamma M)}{2\gamma}(1-1/n)\mathbb{E}_{P_{H}}\big{[}\mathrm{Var}_{\tilde{Z}\sim\mu}(\gamma\ell(H,\tilde{Z}))\big{]}\\ &=\frac{-\exp(-\gamma M)\gamma}{2}(1-1/n)\mathbb{E}_{P_{H}}\big{[}\mathrm{Var}_{\tilde{Z}\sim\mu}(\ell(H,\tilde{Z}))\big{]}\end{split}$ (56)
where $\tilde{Z}\sim\mu$. A similar results also holds for $\gamma<0$. Combining (55), (56) with (54) completes the proof. ∎
We now give a lower bound via the information-theoretic approach.
Theorem 3.14 (Restated).
Under the same assumptions in Theorem 3.13, the expected tilted generalization error satisfies $\overline{\mathrm{gen}}_{\gamma}(H,S)\geq-\frac{(\exp(|\gamma|M)-1)}{|\gamma|}\sqrt{\frac{I(H;S)}{2n}}-\frac{\gamma\exp(\gamma M)}{2}\big{(}1-\frac{1}{n}\big{)}\mathbb{E}_{P_{H}}\big{[}\mathrm{Var}_{\tilde{Z}\sim\mu}(\ell(H,\tilde{Z}))\big{]}.$
Proof.
Similarly as in the proof of Theorem 3.13, we can prove the lower bound. Using the Lipschitz property of the $\log(.)$ function under Assumption 3.1, we have for $\gamma>0$,
$\begin{split}&\mathrm{R}(H,P_{H}\otimes\mu)-\overline{\mathrm{R}}_{\gamma}(H,P_{H}\otimes\mu^{\otimes n})\\ &=\mathbb{E}_{P_{H}\otimes\mu^{\otimes n}}[\frac{1}{\gamma}\log(\exp(\frac{\gamma}{n}\sum_{i=1}^{n}\ell(H,Z_{i})))]-\mathbb{E}_{P_{H}\otimes\mu^{\otimes n}}[\frac{1}{\gamma}\log(\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i})))]\\ &\geq\frac{1}{\gamma}\mathbb{E}_{P_{H}\otimes\mu^{\otimes n}}\Big{[}\exp(\frac{\gamma}{n}\sum_{i=1}^{n}\ell(H,Z_{i}))-\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))\Big{]}\\ &\geq\frac{-\exp(\gamma M)}{2\gamma}\mathbb{E}_{P_{H}\otimes\mu^{\otimes n}}\Bigg{[}\Big{(}\frac{1}{n}\sum_{i=1}^{n}\gamma^{2}\ell(H,Z_{i})^{2}\Big{)}-\Big{(}\frac{1}{n^{2}}\big{(}\sum_{i=1}^{n}\gamma\ell(H,Z_{i})\big{)}^{2}\Big{)}\Bigg{]}\\ &=\frac{-\exp(\gamma M)}{2\gamma}\big{(}1-\frac{1}{n}\big{)}\mathbb{E}_{P_{H}}\big{[}\mathrm{Var}_{\tilde{Z}\sim\mu}(\gamma\ell(H,\tilde{Z}))\big{]}\\ &=\frac{-\gamma\exp(\gamma M)}{2}\big{(}1-\frac{1}{n}\big{)}\mathbb{E}_{P_{H}}\big{[}\mathrm{Var}_{\tilde{Z}\sim\mu}(\ell(H,\tilde{Z}))\big{]}.\end{split}$ (57)
Similar results also holds for $\gamma<0$. Combining (55), (57) with (54) completes the proof. ∎
Corollary 3.15 (Restated).
Under the same assumptions in Theorem 3.13, the absolute value of the expected titled generalization error satisfies $\begin{split}|\overline{\mathrm{gen}}_{\gamma}(H,S)|&\leq\frac{(\exp(|\gamma|M)-1)}{|\gamma|}\sqrt{\frac{I(H;S)}{2n}}+\frac{|\gamma|M^{2}\exp(|\gamma|M)}{8}\Big{(}1-\frac{1}{n}\Big{)}.\end{split}$
Proof.
We can derive the upper bound on absolute value of the expected tilted generalization error by combining Theorem 3.13 and Theorem 3.14. ∎
D.2.1 Individual Sample Bound Discussion
We can apply previous information-theoretic bounding techniques (e.g., (Bu et al., 2020), (Asadi et al., 2018), and (Steinke and Zakynthinou, 2020)), exploiting the Lipschitz property of the logarithm function (or Lemma C.10) over a bounded support. For example, to derive an upper bound based on the individual sample (Bu et al., 2020), using the approach for Proposition 3.12 and Lemma C.10, we have for $\gamma>0$, we have
$\begin{split}&\overline{\mathrm{R}}_{\gamma}(H,P_{H}\otimes\mu^{\otimes n})-\overline{\mathrm{R}}_{\gamma}(H,P_{H,S})\\ &\leq\frac{1}{\gamma}\Big{(}\operatorname{E}_{P_{H}\otimes\mu^{\otimes n}}[\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))]-\operatorname{E}_{P_{H,S}}[\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))]\Big{)}\\ &\leq\frac{\big{(}\exp(\gamma M)-1\big{)}}{\gamma}\sum_{i=1}^{n}\frac{1}{n}\sqrt{\frac{I(H;Z_{i})}{2}},\end{split}$ (58)
where for the first equality, we applied the Lemma C.10 and for the second inequality, we use that $1\leq\exp(\gamma\ell(H,Z_{i}))\leq\exp(\gamma M)$ for $\gamma>0$ and the approach in (Bu et al., 2020) for bounding via individual samples. A similar approach also can be applied to $\gamma<0$. Therefore, the following upper bound holds on the absolute value of the expected generalization error,
$\begin{split}|\overline{\mathrm{gen}}_{\gamma}(H,S)|&\leq\frac{(\exp(|\gamma|M)-1)}{|\gamma|n}\sum_{i=1}^{n}\sqrt{\frac{I(H;Z_{i})}{2n}}+\frac{|\gamma|M^{2}\exp(|\gamma|M)}{8}\Big{(}1-\frac{1}{n}\Big{)}.\end{split}$
D.2.2 Another Approach for an Information-theoretic Bound
Instead of the decomposition (12) we can also consider the following decomposition of the expected tilted generalization error.
$\begin{split}&\mathbb{E}_{P_{H,S}}[\mathrm{gen}_{\gamma}(H,S)]=\mathbb{E}_{P_{H}}[\mathrm{R}(H,\mu)]-\mathrm{R}_{\gamma}(H,P_{H}\otimes\mu)\\ &\quad+\mathrm{R}_{\gamma}(H,P_{H}\otimes\mu)-\mathrm{R}_{\gamma}(H,P_{H,S})+\mathrm{R}_{\gamma}(H,P_{H,S})-\overline{\mathrm{R}}_{\gamma}(H,P_{H,S}).\end{split}$ (59)
Proposition D.9.
Under Assumption 3.1, the following inequality holds,
$\begin{split}&\mathrm{R}_{\gamma}(H,P_{H}\otimes\mu)-\mathrm{R}_{\gamma}(H,P_{H,S})\leq\frac{(\exp(|\gamma|M)-1)}{|\gamma|}\sqrt{\frac{I(H;S)}{2n}}.\end{split}$ (60)
Proof.
For $\gamma<0$ and using the Lipschitz property of the $\log(x)$ function on an interval, we have
$\begin{split}&\mathrm{R}_{\gamma}(H,P_{H}\otimes\mu)-\mathrm{R}_{\gamma}(H,P_{H,S})\\ &=\frac{1}{\gamma}\log(\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))]-\frac{1}{\gamma}\log(\mathbb{E}_{P_{H,S}}[\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))])\\ &\leq\frac{\exp(-\gamma M)}{|\gamma|}\Big{|}\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))]-\mathbb{E}_{P_{H,S}}[\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))]\Big{|}\\ &\leq\frac{(\exp(|\gamma|M)-1)}{|\gamma|}\sqrt{\frac{I(H;S)}{2n}}.\end{split}$ (61)
A similar bound can be derived for $\gamma>0$. ∎
Using Proposition D.9 and Lemma C.3, we can derive the following upper bound on the expected generalization error.
Theorem D.10.
Under Assumption 3.1, the following upper bound holds on the expected tilted generalization error,
$\overline{\mathrm{gen}}_{\gamma}(H,S)\leq\frac{(\exp(|\gamma|M)-1)}{|\gamma|}\sqrt{\frac{I(H;S)}{2n}}+\frac{\mathrm{Var}(\exp(\gamma\ell(H,Z)))}{2\gamma}\Big{(}1-\frac{\exp(-2\gamma M)}{n}\Big{)}.$ (62)
Proof.
We have
$\begin{split}&\mathbb{E}_{H,S}[\mathrm{gen}_{\gamma}(H,S)]\\ &=\mathbb{E}_{H,S}[\mathrm{R}(H,\mu)]-\mathrm{R}_{\gamma}(H,P_{H}\otimes\mu)+\mathrm{R}_{\gamma}(H,P_{H}\otimes\mu)-\mathrm{R}_{\gamma}(H,P_{H,S})\\ &\quad+\mathrm{R}_{\gamma}(H,P_{H,S})-\mathbb{E}_{P_{H,S}}[\widehat{\mathrm{R}}_{\gamma}(H,S)].\end{split}$ (63)
As $0\leq\widehat{\mathrm{R}}_{\gamma}(H,S)\leq M$,
$\mathrm{R}_{\gamma}(H,P_{H,S})-\mathbb{E}_{P_{H,S}}[\widehat{\mathrm{R}}_{\gamma}(H,S)]\leq\frac{\mathrm{Var}(\exp(\gamma\ell(H,Z)))}{2\gamma}.$ (64)
Using Proposition D.9, we have
$\begin{split}|\mathrm{R}_{\gamma}(H,P_{H}\otimes\mu)-\mathrm{R}_{\gamma}(H,P_{H,S})|\leq\frac{(\exp(|\gamma|M)-1)}{|\gamma|}\sqrt{\frac{I(H;S)}{2n}}.\end{split}$ (65)
Under Assumption 3.1, we have
$\begin{split}&\mathbb{E}_{P_{H,S}}[\mathrm{R}(H,\mu)]-\mathrm{R}_{\gamma}(H,P_{H}\otimes\mu)\\ &=\mathbb{E}_{P_{H}\otimes\mu}[\frac{1}{n}\sum_{i=1}^{n}\ell(H,Z_{i})]-\frac{1}{\gamma}\log(\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\frac{1}{n}\sum_{i=1}^{n}\ell(H,Z_{i}))])\\ &\leq-\exp(-2\gamma M)\frac{\mathrm{Var}(\exp(\gamma\ell(H,Z)))}{2\gamma n}.\end{split}$ (66)
Combining (64), (65) and (66) with (63) completes the proof. ∎
Appendix E Proofs and Details of Section 4
E.1 Uniform bounds: details for unbounded loss
Theorem 4.2 (Restated).
Given Assumption 4.1, for any fixed $h\in\mathcal{H}$ with probability at least $(1-\delta)$, then the following upper bound holds on the tilted generalization error for $\gamma<0$, $\begin{split}\mathrm{gen}_{\gamma}(h,S)&\leq 2\kappa_{u}\exp(-\gamma\kappa_{u})\sqrt{\frac{\log(2/\delta)}{n}}-\frac{4\exp(-\gamma\kappa_{u})\log(2/\delta)}{3n\gamma}-\frac{\gamma}{2}\kappa_{u}^{2}.\end{split}$ (67)
Proof.
Using Bernstein’s inequality, Lemma C.7, for $X_{i}=\exp(\gamma\ell(h,Z_{i}))$ and considering $10<X_{i}<$ 1, we have
$\frac{1}{n}\sum_{i=1}^{n}\exp{\gamma\ell(h,Z_{i})}\leq\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))]+\sqrt{\frac{4\mathrm{Var}(\exp(\gamma\ell(h,\tilde{Z})))\log(2/\delta)}{n}}+\frac{4\log(2/\delta)}{3n},$
where we also used that
$\log(x+y)=\log(y)+\log(1+\frac{x}{y})\leq\log(y)+\frac{x}{y}\mbox{ for }y>x>0.$ (68)
Thus,
$\begin{split}&\log\Big{(}\frac{1}{n}\sum_{i=1}^{n}\exp{\gamma\ell(h,Z_{i})}\Big{)}\\ \leq&\log\Big{(}\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))]\Big{)}\\ &+\frac{1}{\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))]}\sqrt{\frac{4\mathrm{Var}(\exp(\gamma\ell(h,\tilde{Z})))\log(2/\delta)}{n}}+\frac{1}{\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))]}\frac{4\log(2/\delta)}{3n}\\ \leq&\log\Big{(}\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))]\Big{)}\\ &+\exp(-\gamma\kappa_{u})\sqrt{\frac{4\mathrm{Var}(\exp(\gamma\ell(h,\tilde{Z})))\log(2/\delta)}{n}}+\exp(-\gamma\kappa_{u})\frac{4\log(2/\delta)}{3n}.\end{split}$
Therefore, for $\gamma<0$ we have
$\begin{split}&\frac{1}{\gamma}\log\Big{(}\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))]\Big{)}-\frac{1}{\gamma}\log\Big{(}\frac{1}{n}\sum_{i=1}^{n}\exp{\gamma\ell(h,Z_{i})}\Big{)}\\ &\leq\frac{\exp(-\gamma\kappa_{u})}{|\gamma|}\sqrt{\frac{4\mathrm{Var}(\exp(\gamma\ell(h,\tilde{Z})))\log(2/\delta)}{n}}+\exp(-\gamma\kappa_{u})\frac{4\log(2/\delta)}{3n|\gamma|}.\end{split}$
Using Lemma C.2, completes the proof. ∎
Theorem 4.3 (Restated).
Given Assumption 4.1, there exists a $\zeta\in(0,1)$ such that for $n\geq\frac{(4\gamma^{2}\kappa_{u}^{2}+8/3\zeta)\log(2/\delta)}{\zeta^{2}\exp(2\gamma\kappa_{u})}$ and $\gamma<0$, for any fixed $h\in\mathcal{H}$, the following lower bound on the tilted generalization error holds with probability at least $(1-\delta)$; $\mathrm{gen}_{\gamma}(h,S)\geq-\frac{2\kappa_{u}\exp(-\gamma\kappa_{u})}{(1-\zeta)}\sqrt{\frac{\log(2/\delta)}{n}}+\frac{4\exp(-\gamma\kappa_{u})(\log(2/\delta))}{3n\gamma(1-\zeta)}.$ (69)
Proof.
Recall that $\tilde{Z}\sim\mu$ and
$\mathrm{gen}_{\gamma}(h,S)$ $=\mathrm{R}(h,\mu)-\frac{1}{\gamma}\log(\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))])$
$\quad+\frac{1}{\gamma}\log(\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))])-\frac{1}{\gamma}\log\left(\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(h,Z_{i}))\right).$
First, we apply Lemma C.9 to yield $\mathrm{R}(h,\mu)-\frac{1}{\gamma}\log(\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))])\geq 0$. Next we focus on the second line of this display. Bernstein’s inequality, Lemma C.7, for $X_{i}=\exp(\gamma\ell(h,Z_{i}))$, so that $10<X_{i}<$ 1, gives that with probability at least $(1-\delta)$,
$\frac{1}{n}\sum_{i=1}^{n}\exp{\gamma\ell(h,Z_{i})}\geq\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))]-\sqrt{\frac{4\operatorname{Var}(\exp(\gamma\ell(h,\tilde{Z})))\log(2/\delta)}{n}}-\frac{4\log(2/\delta)}{3n}.$ (70)
Assume for now that there is a $\zeta\in(0,1)$ such that
$\sqrt{\frac{4\operatorname{Var}(\exp(\gamma\ell(h,\tilde{Z})))\log(2/\delta)}{n}}+\frac{4\log(2/\delta)}{3n}\leq\zeta\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))].$ (71)
As $\log(y-x)=\log(y)+\log(1-\frac{x}{y})\geq\log(y)-\frac{x}{y-x}$ for $y>x>0$, then by taking $y=\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))]$ and $x=\sqrt{\frac{4\operatorname{Var}(\exp(\gamma\ell(h,\tilde{Z})))\log(2/\delta)}{n}}+\frac{4\log(2/\delta)}{3n}$, so that with (71) we have $y-x\geq(1-\zeta)y>0,$ taking logarithms on both sides of (70) gives that with probability at least $(1-\delta),$
$\begin{split}&\log\Big{(}\frac{1}{n}\sum_{i=1}^{n}\exp{\gamma\ell(h,Z_{i})}\Big{)}\\ &\geq\log\Big{(}\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))]\Big{)}-\frac{1}{\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))]}\left(\sqrt{\frac{4\operatorname{Var}(\exp(\gamma\ell(h,\tilde{Z})))\log(2/\delta)}{n}}+\frac{4\log(2/\delta)}{3n}\right)\\ &\geq\log\Big{(}\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))]\Big{)}-\frac{1}{(1-\zeta)\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))]}\sqrt{\frac{4\operatorname{Var}(\exp(\gamma\ell(h,\tilde{Z})))\log(2/\delta)}{n}}\\ &\qquad-\frac{1}{(1-\zeta)\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))]}\frac{4\log(2/\delta)}{3n}\\ &\geq\log\Big{(}\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))]\Big{)}-\frac{\exp(-\gamma\kappa_{u})}{(1-\zeta)}\sqrt{\frac{4\operatorname{Var}(\exp(\gamma\ell(h,\tilde{Z})))\log(2/\delta)}{n}}-\frac{\exp(-\gamma\kappa_{u})}{(1-\zeta)}\frac{4\log(2/\delta)}{3n}\\ &\geq\log\Big{(}\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))]\Big{)}-\frac{|\gamma|2\kappa_{u}\exp(-\gamma\kappa_{u})}{(1-\zeta)}\sqrt{\frac{\log(2/\delta)}{n}}-\frac{\exp(-\gamma\kappa_{u})}{(1-\zeta)}\frac{4\log(2/\delta)}{3n}.\end{split}$ (72)
Here we used that by Assumption 4.1 and Lemma C.8,
$\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))]\leq\exp[\mathbb{E}(\gamma\ell(h,\tilde{Z}))]\leq\exp(\gamma\kappa_{u})$
and by Assumption 4.1,
$\mathrm{Var}(\exp(\gamma\ell(h,\tilde{Z})))\leq[\mathbb{E}{\exp(\gamma\ell(h,\tilde{Z}))^{2}}]\leq\exp[\mathbb{E}{(\gamma\ell(h,\tilde{Z}))^{2}}]\leq\exp(\gamma^{2}\kappa_{u}^{2}).$
This gives the stated bound assuming that (71) holds. In order to satisfy (71), viewing (71) as a quadratic inequality in $\sqrt{n}$ and using that $(a+b)^{2}\leq 2a^{2}+2b^{2}$ yields
$n\geq\frac{(4\operatorname{Var}(\exp(\gamma\ell(h,\tilde{Z})))+8/3\zeta\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))])\log(2/\delta)}{\zeta^{2}(\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))])^{2}},$
Now applying $\exp(\gamma\kappa_{u})\leq\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))]\leq 1$ and $\operatorname{Var}(\exp(\gamma\ell(h,\tilde{Z})))\leq\gamma^{2}\kappa_{u}^{2}$ completes the proof. ∎
Corollary 4.4 (Restated).
Under the same assumptions in Theorem 4.3 and a finite hypothesis space, then for $n\geq\frac{(4\gamma^{2}\kappa_{u}^{2}+8/3\zeta)\log(2/\delta)}{\zeta^{2}\exp(2\gamma\kappa_{u})}$, for $\gamma<0$ and with probability at least $(1-\delta)$, the absolute value of the titled generalization error satisfies $\begin{split}\sup_{h\in\mathcal{H}}|\mathrm{gen}_{\gamma}(h,S)|&\leq\frac{2\kappa_{u}\exp(-\gamma\kappa_{u})}{(1-\zeta)}\sqrt{\frac{(\log(\mathrm{card}(\mathcal{H}))+\log(2/\delta))}{n}}\\ &\qquad-\frac{4\exp(-\gamma\kappa_{u})(\log(\mathrm{card}(\mathcal{H}))+\log(2/\delta))}{3n\gamma(1-\zeta)}-\frac{\gamma}{2}\kappa_{u}^{2}.\end{split}$ (73)
Proof.
Combining the upper and lower bounds, Theorem 4.2 and Theorem 4.3, we can derive the following bound for a fixed $h\in\mathcal{H}$,
$\begin{split}|\mathrm{gen}_{\gamma}(h,S)|&\leq\frac{2\kappa_{u}\exp(-\gamma\kappa_{u})}{(1-\zeta)}\sqrt{\frac{(\log(2/\delta))}{n}}\\ &\qquad-\frac{4\exp(-\gamma\kappa_{u})(\log(2/\delta))}{3n\gamma(1-\zeta)}-\frac{\gamma}{2}\kappa_{u}^{2}.\end{split}$ (74)
Then, using the uniform bound Lemma C.11 completes the proof. ∎
E.2 Information-theoretic Bounds
In the information-theoretic approach for the unbounded loss function, we relax the uniform assumption, Assumption 4.1, as follows,
Assumption E.1 (Bounded second moment).
The learning algorithm $P_{H|S}$, loss function $\ell$, and $\mu$ are such that there is a constant $\kappa_{t}\in\mathbb{R}^{+}$ where the loss function $(H,Z)\mapsto\ell(H,Z)$ satisfies $\mathbb{E}_{P_{H}\otimes\mu}[\ell^{2}(H,Z)]\leq\kappa_{t}^{2}$.
The following results are helpful in deriving the upper and lower bounds under unbounded loss assumptions via the information-theoretic approach.
Proposition E.2.
Given Assumption E.1, the following inequality holds for $\gamma<0$,
$\begin{split}\mathrm{R}_{\gamma}(H,P_{H}\otimes\mu)-\mathrm{R}_{\gamma}(H,P_{H,S})\leq\begin{cases}\exp(-\gamma\kappa_{t})\sqrt{\frac{\kappa_{t}^{2}I(H;S)}{n}},\qquad&\mbox{ if }\frac{I(H;S)}{n}\leq\frac{\gamma^{2}\kappa_{t}^{2}}{2}\\ \frac{-\exp(-\gamma\kappa_{t})}{\gamma}\Big{(}\frac{I(H;S)}{n}+\frac{\gamma^{2}\kappa_{t}^{2}}{2}\Big{)},\qquad&\mbox{ if }\frac{I(H;S)}{n}>\frac{\gamma^{2}\kappa_{t}^{2}}{2}\end{cases}\end{split}$
Proof.
For $\gamma<0$ and we use that $0\leq\exp(\gamma\ell(H,\tilde{Z}))\leq 1$ and $\mathrm{Var}(\exp(\gamma\ell(H,\tilde{Z})))\leq\gamma^{2}\mathrm{Var}(\ell(H,\tilde{Z}))\leq\gamma^{2}\kappa_{t}^{2}$. Note that the variable $\exp(\gamma\ell(H,\tilde{Z}))$ is sub-exponential with parameters $(\gamma^{2}\kappa_{t}^{2},1)$ under the distribution $P_{H}\otimes\mu$. Using the approach in (Bu et al., 2020; Aminian et al., 2021a) for the sub-exponential case, we have
$\begin{split}&\Big{|}\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))]-\mathbb{E}_{P_{H,S}}[\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))]\Big{|}\\ &\leq\begin{cases}|\gamma|\sqrt{\kappa_{t}^{2}\frac{I(H;S)}{n}}&\quad\mbox{ if}\frac{I(H;S)}{n}\leq\frac{\gamma^{2}\kappa_{t}^{2}}{2}\\ \frac{I(H;S)}{n}+\frac{\gamma^{2}\kappa_{t}^{2}}{2}&\mbox{ if }\frac{I(H;S)}{n}>\frac{\gamma^{2}\kappa_{t}^{2}}{2}.\end{cases}\end{split}$ (75)
Therefore, we have for $\frac{I(H;S)}{n}\leq\frac{\gamma^{2}\kappa_{t}^{2}}{2}$,
$\begin{split}\mathbb{E}_{P_{H,S}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))\Big{]}\leq\Big{(}\mathbb{E}_{P_{H}\otimes\mu}\Big{[}\exp(\gamma\ell(H,\tilde{Z}))\Big{]}+|\gamma|\sqrt{\frac{\kappa_{t}^{2}I(H;S)}{n}}\Big{)}.\end{split}$ (76)
Using (68) gives
$\begin{split}&\frac{1}{\gamma}\log\Big{(}\mathbb{E}_{P_{H,S}}[\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))]\Big{)}-\frac{1}{\gamma}\log(\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))])\\ &\geq\frac{|\gamma|}{\gamma\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))]}\sqrt{\frac{\kappa_{t}^{2}I(H;S)}{n}}.\end{split}$ (77)
For $\frac{I(H;S)}{n}>\frac{\gamma^{2}\kappa_{t}^{2}}{2}$, we have
$\begin{split}&\mathbb{E}_{P_{H,S}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))\Big{]}\\ &\leq\Big{(}\mathbb{E}_{P_{H}\otimes\mu}\Big{[}\exp(\gamma\ell(H,\tilde{Z}))\Big{]}+\frac{I(H;S)}{n}+\frac{\gamma^{2}\kappa_{t}^{2}}{2}\Big{)}.\end{split}$ (78)
Using (68) again, we obtain,
$\begin{split}&\frac{1}{\gamma}\log\Big{(}\mathbb{E}_{P_{H,S}}[\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))]\Big{)}-\frac{1}{\gamma}\log(\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))])\\ &\geq\frac{1}{\gamma\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))]}\Big{(}\frac{I(H;S)}{n}+\frac{\gamma^{2}\kappa_{t}^{2}}{2}\Big{)}.\end{split}$ (79)
As under Assumption E.1 and Lemma C.8, we have $\exp(\gamma\kappa_{t})\leq\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))]$, the final result follows. ∎
Corollary E.3.
Given Assumption E.1, the following inequality holds for $\gamma<0$ and some $\zeta\in(0,1)$,
$\begin{split}\mathrm{R}_{\gamma}(H,P_{H}\otimes\mu)-\mathrm{R}_{\gamma}(H,P_{H,S})\geq\begin{cases}\frac{-\exp(-\gamma\kappa_{t})}{(1-\zeta)}\sqrt{\frac{\kappa_{t}^{2}I(H;S)}{n}},\quad&\mbox{ if }\max\big{(}\frac{2I(H;S)}{\gamma^{2}\kappa_{t}^{2}},\frac{\gamma^{2}\kappa_{t}^{2}I(H;S)}{\zeta^{2}\exp(2\gamma\kappa_{t})}\big{)}\leq n\\ \frac{\exp(-\gamma\kappa_{t})}{(1-\zeta)\gamma}\Big{(}\frac{I(H;S)}{n}+\frac{\gamma^{2}\kappa_{t}^{2}}{2}\Big{)},\quad&\mbox{ if }\min(\frac{4I(H;S)\big{(}4\zeta^{-2}-1\big{)}}{\gamma^{2}\kappa_{t}^{2}},\frac{2I(H;S)}{\gamma^{2}\kappa_{t}^{2}})>n.\end{cases}\end{split}$
Proof.
Similar to Proposition E.2, we have
$\begin{split}&\Big{|}\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))]-\mathbb{E}_{P_{H,S}}[\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))]\Big{|}\\ &\leq\begin{cases}|\gamma|\sqrt{\kappa_{t}^{2}\frac{I(H;S)}{n}}&\quad\mbox{ if}\frac{I(H;S)}{n}\leq\frac{\gamma^{2}\kappa_{t}^{2}}{2}\\ \frac{I(H;S)}{n}+\frac{\gamma^{2}\kappa_{t}^{2}}{2}&\mbox{ if }\frac{I(H;S)}{n}>\frac{\gamma^{2}\kappa_{t}^{2}}{2}.\end{cases}\end{split}$ (80)
Therefore, we have for $\frac{I(H;S)}{n}\leq\frac{\gamma^{2}\kappa_{t}^{2}}{2}$,
$\begin{split}\mathbb{E}_{P_{H,S}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))\Big{]}\geq\Big{(}\mathbb{E}_{P_{H}\otimes\mu}\Big{[}\exp(\gamma\ell(H,\tilde{Z}))\Big{]}+\gamma\sqrt{\frac{\kappa_{t}^{2}I(H;S)}{n}}\Big{)}.\end{split}$ (81)
Assume for now that there is a $\zeta\in(0,1)$ such that
$|\gamma|\sqrt{\frac{\kappa_{t}^{2}I(H;S)}{n}}\leq\zeta\mathbb{E}_{P_{H}\otimes\mu}\Big{[}\exp(\gamma\ell(H,\tilde{Z}))\Big{]}$ (82)
As $\log(y-x)=\log(y)+\log(1-\frac{x}{y})\geq\log(y)-\frac{x}{y-x}$ for $y>x>0$, then by taking $y=\mathbb{E}_{P_{H}\otimes\mu}\Big{[}\exp(\gamma\ell(H,\tilde{Z}))\Big{]}$ and $x=|\gamma|\sqrt{\frac{\kappa_{t}^{2}I(H;S)}{n}}$, so that with (71) we have $y-x\geq(1-\zeta)y>0,$ taking logarithms on both sides of (81) gives that
$\begin{split}&\frac{1}{\gamma}\log\Big{(}\mathbb{E}_{P_{H,S}}[\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))]\Big{)}-\frac{1}{\gamma}\log(\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))])\\ &\geq\frac{1}{(1-\zeta)\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))]}\sqrt{\frac{\kappa_{t}^{2}I(H;S)}{n}},\end{split}$ (83)
where it holds for $n\geq\frac{\gamma^{2}\kappa_{t}^{2}I(H;S)}{\zeta^{2}\exp(2\gamma\kappa_{t})}$. For $\frac{I(H;S)}{n}>\frac{\gamma^{2}\kappa_{t}^{2}}{2}$, we have
$\begin{split}&\mathbb{E}_{P_{H,S}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))\Big{]}\geq\Big{(}\mathbb{E}_{P_{H}\otimes\mu}\Big{[}\exp(\gamma\ell(H,\tilde{Z}))\Big{]}-\frac{I(H;S)}{n}-\frac{\gamma^{2}\kappa_{t}^{2}}{2}\Big{)}.\end{split}$ (84)
As $\log(y-x)=\log(y)+\log(1-\frac{x}{y})\geq\log(y)-\frac{x}{y-x}$, we obtain,
$\begin{split}&\frac{1}{\gamma}\log\Big{(}\mathbb{E}_{P_{H,S}}[\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))]\Big{)}-\frac{1}{\gamma}\log(\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))])\\ &\geq\frac{1}{\gamma(1-\zeta)\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))]}\Big{(}\frac{I(H;S)}{n}+\frac{\gamma^{2}\kappa_{t}^{2}}{2}\Big{)},\end{split}$ (85)
where it holds for $\frac{4I(H;S)\big{(}4\zeta^{-2}-1\big{)}}{\gamma^{2}\kappa_{t}^{2}}\geq n$. Note that, we have,
$\begin{split}&\mathbb{E}_{P_{H}\otimes\mu}\Big{[}\exp(\gamma\ell(H,\tilde{Z}))\Big{]}\\ &\geq\frac{|\gamma|}{\zeta}\sqrt{\frac{\kappa_{t}^{2}I(H;S)}{n}}\\ &\geq\frac{I(H;S)}{n}+\frac{\gamma^{2}\kappa_{t}^{2}}{2}.\end{split}$ (86)
Therefore, we have,
$\frac{|\gamma|}{\zeta}\sqrt{\frac{\kappa_{t}^{2}I(H;S)}{n}}\geq\frac{I(H;S)}{n}+\frac{\gamma^{2}\kappa_{t}^{2}}{2},$ (87)
where 87 is a quadratic inequality in $\sqrt{n}$ and using that $(a+b)^{2}\leq 2a^{2}+2b^{2}$ yields,
$\frac{4I(H;S)\big{(}4\zeta^{-2}-1\big{)}}{\gamma^{2}\kappa_{t}^{2}}\geq n.$
As under Assumption E.1 and Lemma C.8, we have $\exp(\gamma\kappa_{t})\leq\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))]$, the final result follows. ∎
Using Proposition E.2, we can obtain the following upper bound on the expected generalization error.
Theorem E.4.
Given Assumption E.1, the following upper bound holds on the expected tilted generalization error for $\gamma<0$,
$\begin{split}\overline{\mathrm{gen}}_{\gamma}(H,S)&\leq\begin{cases}\exp(-\gamma\kappa_{t})\sqrt{\frac{\kappa_{t}^{2}I(H;S)}{n}}-\frac{\gamma}{2}\kappa_{t}^{2},\qquad&\mbox{ if }\frac{I(H;S)}{n}\leq\frac{\gamma^{2}\kappa_{t}^{2}}{2}\\ \frac{-\exp(-\gamma\kappa_{t})}{\gamma}\Big{(}\frac{I(H;S)}{n}+\frac{\gamma^{2}\kappa_{t}^{2}}{2}\Big{)}-\frac{\gamma}{2}\kappa_{t}^{2},\qquad&\mbox{ if }\frac{I(H;S)}{n}>\frac{\gamma^{2}\kappa_{t}^{2}}{2}\end{cases}\end{split}$
Proof.
We use the following decomposition,
$\begin{split}&\mathbb{E}_{P_{H,S}}[\mathrm{gen}_{\gamma}(H,S)]\\ &=\mathbb{E}_{H,S}[\mathrm{R}(H,\mu)]-\mathrm{R}_{\gamma}(H,P_{H}\otimes\mu)+\mathrm{R}_{\gamma}(H,P_{H}\otimes\mu)-\mathrm{R}_{\gamma}(H,P_{H,S})\\ &\quad+\mathrm{R}_{\gamma}(H,P_{H,S})-\mathbb{E}_{P_{H,S}}[\widehat{\mathrm{R}}_{\gamma}(H,S)].\end{split}$ (88)
Using Lemma C.9, we have
$\begin{split}&\mathbb{E}_{P_{H,S}}[\mathrm{R}(H,\mu)]-\mathrm{R}_{\gamma}(H,P_{H}\otimes\mu)\leq\frac{|\gamma|}{2}\mathbb{E}_{P_{H}\otimes\mu}[\ell^{2}(H,\tilde{Z})],\end{split}$ (89)
and due to Jensen’s inequality for $\gamma<0$, we have
$\mathrm{R}_{\gamma}(H,P_{H,S})-\mathbb{E}_{P_{H,S}}[\widehat{\mathrm{R}}_{\gamma}(H,S)]\leq 0.$ (90)
Using a similar approach as for the proof of Proposition D.9 and applying Proposition E.2, we obtain
$\begin{split}\mathrm{R}_{\gamma}(H,P_{H}\otimes\mu)-\mathrm{R}_{\gamma}(H,P_{H,S})\leq\begin{cases}\exp(-\gamma\kappa_{t})\sqrt{\frac{\kappa_{t}^{2}I(H;S)}{n}},\qquad&\frac{I(H;S)}{n}\leq\frac{\gamma^{2}\kappa_{t}^{2}}{2}\\ \frac{\exp(-\gamma\kappa_{t})}{|\gamma|}\Big{(}\frac{I(H;S)}{n}+\frac{\gamma^{2}\kappa_{t}^{2}}{2}\Big{)},\qquad&\frac{I(H;S)}{n}>\frac{\gamma^{2}\kappa_{t}^{2}}{2}.\end{cases}\end{split}$ (91)
Combining (90), (91) and (89) with (88) completes the proof. ∎
Remark E.5.
Assuming $\gamma=O(n^{-1/2})$, the upper bound in Theorem E.4 has the convergence rate $O(n^{-1/2})$. Note that the result in Theorem E.4 holds for unbounded loss functions, provided that the second moment of the loss function exists.
A similar approach to Theorem E.4 can be applied to derive an information-theoretical lower bound on the expected tilted generalization error for $\gamma<0$ by applying Corollary E.3.
Theorem E.6.
Given Assumption E.1, the following lower bound holds on the expected tilted generalization error for $\gamma<0$,
$\begin{split}\overline{\mathrm{gen}}_{\gamma}(H,S)&\geq\begin{cases}-\exp(-\gamma\kappa_{t})\sqrt{\frac{\kappa_{t}^{2}I(H;S)}{n}}+\frac{\gamma}{2}\kappa_{t}^{2},\qquad&\mbox{ if }\max\big{(}\frac{2I(H;S)}{\gamma^{2}\kappa_{t}^{2}},\frac{\gamma^{2}\kappa_{t}^{2}I(H;S)}{\zeta^{2}\exp(2\gamma\kappa_{t})}\big{)}\leq n\\ \frac{\exp(-\gamma\kappa_{t})}{\gamma}\Big{(}\frac{I(H;S)}{n}+\frac{\gamma^{2}\kappa_{t}^{2}}{2}\Big{)}+\frac{\gamma}{2}\kappa_{t}^{2},\qquad&\mbox{ if }\min(\frac{4\big{(}4\zeta^{-2}-1\big{)}I(H;S)}{\gamma^{2}\kappa_{t}^{2}},\frac{2I(H;S)}{\gamma^{2}\kappa_{t}^{2}})>n.\end{cases}\end{split}$
Proof.
The proof is similar to Theorem E.4 and using Corollary E.3. ∎
Combining Theorem E.6 and Theorem E.4, we can derive the upper bound on the absolute value of the expected generalization error.
Appendix F Proof and details of Section 5
Proposition 5.2 (Restated).
Under Assumption 4.1 and Assumption 5.1, the difference of the tilted population risk (9) between $\mu$ and $\tilde{\mu}$ is bounded as follows;
$\begin{split}&\Big{|}\frac{1}{\gamma}\log(\mathbb{E}_{\tilde{Z}\sim\mu}[\exp(\gamma\ell(h,\tilde{Z}))])-\frac{1}{\gamma}\log(\mathbb{E}_{\tilde{Z}\sim\tilde{\mu}}[\exp(\gamma\ell(h,\tilde{Z}))])\Big{|}\leq\frac{\mathbb{TV}(\mu,\tilde{\mu})}{\gamma^{2}}\frac{\exp(|\gamma|\kappa_{u})-\exp(|\gamma|\kappa_{s})}{(\kappa_{u}-\kappa_{s})}.\end{split}$ (92)
Proof.
We have for a fixed $h\in\mathcal{H}$ that
$\begin{split}&\Big{|}\frac{1}{\gamma}\log(\mathbb{E}_{\tilde{Z}\sim\mu}[\exp(\gamma\ell(h,\tilde{Z}))])-\frac{1}{\gamma}\log(\mathbb{E}_{\tilde{Z}\sim\tilde{\mu}}[\exp(\gamma\ell(h,\tilde{Z}))])\Big{|}\\ &\overset{(a)}{=}\Big{|}\int_{0}^{1}\int_{\mathcal{Z}}\frac{\exp(\gamma\ell(h,z))}{|\gamma|\mathbb{E}_{\tilde{Z}\sim\mu_{\lambda}}[\exp(\gamma\ell(h,\tilde{Z}))]}(\tilde{\mu}-\mu)(\mathrm{d}z)\mathrm{d}\lambda\Big{|}\\ &\overset{(b)}{\leq}\frac{\mathbb{TV}(\mu,\tilde{\mu})\exp(|\gamma|\kappa_{s})}{|\gamma|}\int_{0}^{1}\exp(|\gamma|\lambda(\kappa_{u}-\kappa_{s}))\mathrm{d}\lambda\\ &=\frac{\mathbb{TV}(\mu,\tilde{\mu})}{|\gamma|}\frac{\exp(|\gamma|\kappa_{u})-\exp(|\gamma|\kappa_{s})}{|\gamma|(\kappa_{u}-\kappa_{s})},\end{split}$ (93)
where (a) and (b) follow from the functional derivative with $\mu_{\lambda}=\tilde{\mu}+\lambda(\mu-\tilde{\mu})$ and Lemma C.16. The same approach can be applied for lower bound. ∎
Proposition 5.3 (Restated).
Given Assumptions  4.1 and 5.1, for any fixed $h\in\mathcal{H}$ and with probability least $(1-\delta)$ for $\gamma<0$, then the following upper bound holds on the tilted generalization error $\begin{split}\mathrm{gen}_{\gamma}(h,\hat{S})&\leq 2\kappa_{s}\exp(-\gamma\kappa_{s})\sqrt{\frac{(\log(2/\delta))}{n}}\\ &\qquad-\frac{4\exp(-\gamma\kappa_{s})(\log(2/\delta))}{3n\gamma}-\frac{\gamma}{2}\kappa_{u}^{2}+\frac{\mathbb{TV}(\mu,\tilde{\mu})}{\gamma^{2}}\frac{\big{(}\exp(|\gamma|\kappa_{u})-\exp(|\gamma|\kappa_{s})\big{)}}{(\kappa_{u}-\kappa_{s})},\end{split}$ where $\hat{S}$ is the training dataset under the distributional shift.
Proof.
The proof follows directly from the following decomposition of the tilted generalization error under distribution shift,
$\mathrm{gen}_{\gamma}(h,\hat{S})=\underbrace{\mathrm{R}(h,\mu)-\mathrm{R}_{\gamma}(h,\mu^{\otimes n})}_{I_{5}}+\underbrace{\mathrm{R}_{\gamma}(h,\mu)-\mathrm{R}_{\gamma}(h,\tilde{\mu})}_{I_{6}}+\underbrace{\mathrm{R}_{\gamma}(h,\tilde{\mu})-\widehat{\mathrm{R}}_{\gamma}(h,\hat{S})}_{I_{7}},$
where $I_{5}$, $I_{6}$ and $I_{7}$ can be bounded using Lemma C.9, Proposition 5.2 and Theorem 4.2, respectively. ∎
Remark F.1.
It is noteworthy that the upper bound in Proposition 5.3 can be infinite for $\gamma\rightarrow-\infty$ and $\gamma=0$ and a fixed n. Consequently, there must exist a $\gamma\in(-\infty,0)$ that minimizes this upper bound. To illustrate this point, consider the case where $n\rightarrow\infty$; here, the first and second terms in the upper bound would vanish. Thus, we are led to the following minimization problem:
$\gamma^{\star}:=\mathop{\mathrm{arg}\min}_{\gamma\in(-\infty,0)}\left[-\frac{\gamma}{2}\kappa_{u}^{2}+\frac{\mathbb{TV}(\mu,\tilde{\mu})}{\gamma^{2}}\frac{\big{(}\exp(|\gamma|\kappa_{u})-\exp(|\gamma|\kappa_{s})\big{)}}{(\kappa_{u}-\kappa_{s})}\right],$ (94)
where the solution exists. As $\gamma^{\star}$ decreases when $\mathbb{TV}(\mu,\tilde{\mu})$ increases, practically, this implies that if the training distribution becomes more adversarial (i.e., further away from the benign test distribution), we would use smaller negative $\gamma$’s to bypass outliers.
Proposition 5.4 (Restated).
Given Assumptions  4.1 and 5.1, for any fixed $h\in\mathcal{H}$ and with probability least $(1-\delta)$, there exists a $\zeta\in(0,1)$ such that for $n\geq\frac{(4\gamma^{2}\kappa_{u}^{2}+8/3\zeta)\log(2/\delta)}{\zeta^{2}\exp(2\gamma\kappa_{u})}$ and $\gamma<0$, such that the following upper bound holds on the tilted generalization error $\begin{split}\mathrm{gen}_{\gamma}(h,\hat{S})&\geq-\frac{2\kappa_{s}\exp(-\gamma\kappa_{s})}{(1-\zeta)}\sqrt{\frac{\log(2/\delta)}{n}}+\frac{4\exp(-\gamma\kappa_{s})(\log(2/\delta))}{3n\gamma(1-\zeta)}\\ &\qquad-\frac{\mathbb{TV}(\mu,\tilde{\mu})}{\gamma^{2}}\frac{\big{(}\exp(|\gamma|\kappa_{u})-\exp(|\gamma|\kappa_{s})\big{)}}{(\kappa_{u}-\kappa_{s})},\end{split}$ where $\hat{S}$ is the training dataset under the distributional shift.
Proof.
The proof follows directly from the following decomposition of the tilted generalization error under distribution shift,
$\mathrm{gen}_{\gamma}(h,\hat{S})=\underbrace{\mathrm{R}(h,\mu)-\mathrm{R}_{\gamma}(h,\mu^{\otimes n})}_{I_{5}}+\underbrace{\mathrm{R}_{\gamma}(h,\mu)-\mathrm{R}_{\gamma}(h,\tilde{\mu})}_{I_{6}}+\underbrace{\mathrm{R}_{\gamma}(h,\tilde{\mu})-\widehat{\mathrm{R}}_{\gamma}(h,\hat{S})}_{I_{7}},$
where $I_{5}$, $I_{6}$ and $I_{7}$ can be bounded using Lemma C.9, Proposition 5.2 and Theorem 4.3, respectively. ∎
Theorem 5.5 (Restated).
Under the same assumptions in Proposition 5.4, then for $n\geq\frac{(4\gamma^{2}\kappa_{u}^{2}+8/3\zeta)\log(2/\delta)}{\zeta^{2}\exp(2\gamma\kappa_{u})}$ and $\gamma<0$, and with probability at least $(1-\delta)$, the absolute value of the tilted generalization error under distributional shift satisfies $\begin{split}\sup_{h\in\mathcal{H}}|\mathrm{gen}_{\gamma}(h,\hat{S})|&\leq\frac{2\kappa_{s}\exp(-\gamma\kappa_{s})}{(1-\zeta)}\sqrt{\frac{B(\delta)}{n}}-\frac{4\exp(-\gamma\kappa_{s})B(\delta)}{3n\gamma(1-\zeta)}\\ &\qquad-\frac{\gamma\kappa_{u}^{2}}{2}+\frac{\mathbb{TV}(\mu,\tilde{\mu})}{\gamma^{2}}\frac{\big{(}\exp(|\gamma|\kappa_{u})-\exp(|\gamma|\kappa_{s})\big{)}}{(\kappa_{u}-\kappa_{s})},\end{split}$ (95) where $B(\delta)=\log(\mathrm{card}(\mathcal{H}))+\log(2/\delta)$.
Proof.
The result follows from combining the results of Proposition 5.3, Proposition 5.4 and applying the uniform bound. ∎
Appendix G Proofs and details of Section 6
Proposition 6.1 (Restated).
The solution to the expected TERM regularized via KL divergence, (26), is the tilted Gibbs posterior,
$P_{H|S}^{\gamma}=\frac{\pi_{H}}{F_{\alpha}(S)}\Big{(}\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,z_{i}))\Big{)}^{-\alpha/\gamma},$ (96)
where $F_{\alpha}(S)$ is the normalization factor.
Proof.
From (Zhang, 2006), we know that,
$P_{X}^{\star}=\min_{P_{X}}\mathbb{E}_{P_{X}}[f(x)]+\frac{1}{\alpha}\mathrm{KL}(P_{X}\|Q_{X}),$ (97)
where $P_{X}^{\star}=\frac{Q_{X}\exp(-\alpha f(X))}{\mathbb{E}_{Q_{X}}[\exp(-\alpha f(X))]}.$ Using (97), it can be shown that the tilted Gibbs posterior is the solution to (96).
Proposition 6.2 (Restated).
The difference between the expected TER under the joint and product of marginal distributions of $H$ and $S$ can be characterized as, $\begin{split}&\overline{\mathrm{R}}_{\gamma}(H,P_{H}\otimes\mu)-\overline{\mathrm{R}}_{\gamma}(H,P_{H,S})=\frac{I_{\mathrm{SKL}}(H;S)}{\alpha}.\end{split}$ (98)
Proof.
As in Aminian et al. (2015), the symmetrized KL information between two random variables $(S,H)$ can be written as
$I_{\mathrm{SKL}}(H;S)=\mathbb{E}_{P_{H}\otimes\mu^{\otimes n}}[\log(P_{H|S})]-\mathbb{E}_{P_{H,S}}[\log(P_{H|S})].$ (99)
The results follows by substituting the tilted Gibbs posterior in (99). ∎
Theorem 6.3 (Restated).
Under Assumption 3.1, the expected generalization error of the tilted Gibbs posterior satisfies, $\overline{\mathrm{gen}}_{\gamma}(H,S)\leq\frac{\alpha(\exp(|\gamma|M)-1)^{2}}{2\gamma^{2}n}+\frac{\mathrm{Var}(\exp(\gamma\ell(H,\tilde{Z})))}{2\gamma}\Big{(}1/n-\exp(-2\gamma M)\Big{)}.$ (100)
Proof.
Note that, we have
$\begin{split}\frac{I(H;S)}{\alpha}&\leq\frac{I_{\mathrm{SKL}}(H;S)}{\alpha}\\ &=\overline{\mathrm{R}}_{\gamma}(H,P_{H}\otimes\mu)-\overline{\mathrm{R}}_{\gamma}(H,P_{H,S})\\ &\leq\Big{|}\overline{\mathrm{R}}_{\gamma}(H,P_{H}\otimes\mu^{\otimes n})-\overline{\mathrm{R}}_{\gamma}(H,P_{H,S})\Big{|}\\ &\leq\frac{(\exp(|\gamma|M)-1)}{|\gamma|}\sqrt{\frac{I(H;S)}{2n}}.\end{split}$ (101)
Therefore, we have
$\frac{I(H;S)}{\alpha}\leq\frac{(\exp(|\gamma|M)-1)}{|\gamma|}\sqrt{\frac{I(H;S)}{2n}}.$ (102)
Solving (102), results in,
$\sqrt{I(H;S)}\leq\alpha\frac{(\exp(|\gamma|M)-1)}{|\gamma|}\sqrt{\frac{1}{2n}}.$ (103)
Therefore, we obtain,
$\begin{split}&\Big{|}\overline{\mathrm{R}}_{\gamma}(H,P_{H}\otimes\mu^{\otimes n})-\overline{\mathrm{R}}_{\gamma}(H,P_{H,S})\Big{|}\\ &\leq\alpha\frac{(\exp(|\gamma|M)-1)^{2}}{2\gamma^{2}n}.\end{split}$ (104)
Using Theorem 3.13, the final result follows. ∎
In addition to KL-regularized linear risk minimization, the Gibbs posterior is also the solution to another problem. For this formulation we recall that the $\alpha$-Rényi divergence between $P$ and $Q$ is given by $\mathrm{R}_{\alpha}(P\|Q):=\frac{1}{\alpha-1}\log\Big{(}\int_{\mathcal{X}}\Big{(}\frac{\mathrm{d}P}{\mathrm{d}Q}\Big{)}^{\alpha}\mathrm{d}Q\Big{)},$ for $\alpha\in(0,1)\cup(1,\infty)$. We also define the conditional Rényi divergence between $P_{X|Y}$ and $Q_{X|Y}$ as $\mathrm{R}_{\alpha}(P_{X|Y}\|Q_{X|Y}|P_{Y}):=\frac{1}{\alpha-1}\mathbb{E}_{P_{Y}}\Big{[}\log\Big{(}\int_{\mathcal{X}}\Big{(}\frac{\mathrm{d}P_{X|Y}}{\mathrm{d}Q_{X|Y}}\Big{)}^{\alpha}\mathrm{d}Q_{X|Y}\Big{)}\Big{]},$ for $\alpha\in(0,1)\cup(1,\infty)$. Here, $P_{X|Y}$ denotes the conditional distribution of $X$ given $Y$.
Proposition G.1 (Gibbs posterior).
Suppose that $\gamma=\frac{1}{\alpha}-1$ and $\alpha\in(0,1)\cup(1,\infty)$. Then the solution to the minimization problem
$P_{H|S}^{\alpha}=\mathop{\mathrm{arg}\,\inf}_{P_{H|S}}\left\{\mathbb{E}_{P_{S}}\Big{[}\frac{1}{\gamma}\log\Big{(}\mathbb{E}_{P_{H|S}}\Big{[}\exp\Big{(}\gamma\hat{\mathrm{R}}(H,S)\Big{)}\Big{]}\Big{)}\Big{]}+\mathrm{R}_{\alpha}(P_{H|S}\|\pi_{H}|P_{S})\right\},$ (105)
with $\hat{\mathrm{R}}(H,S)$ the linear empirical risk (1), is the Gibbs posterior,
$P_{H|S}^{\alpha}=\frac{\pi_{H}[\exp(-\gamma\hat{\mathrm{R}}(H,S))]}{\mathbb{E}_{\pi_{H}}[\exp(-\gamma\hat{\mathrm{R}}(H,S))]},$
where $\pi_{H}$ is the prior distribution on the space $\mathcal{H}$ of hypotheses.
Proof.
Let us consider the following minimization problem,
[…]
where $\gamma=\frac{1}{\alpha}-1$. As shown by Dvijotham and Todorov (2012), the solution to (106) is the Gibbs posterior,
$P_{Y}^{\star}=\frac{Q_{Y}\exp(-\alpha f(Y))}{\mathbb{E}_{Q_{Y}}[\exp(-\alpha f(Y))]}.$
If $\alpha\rightarrow 1$, then $\gamma\rightarrow 0$ and (105) converges to the KL-regularized ERM problem.
The tilted generalization error under the Gibbs posterior can be bounded as follows.
Proposition G.2.
Under Assumption 3.1 when training with the Gibbs posterior, (28), the following upper bound holds on the expected tilted generalization error,
$\overline{\mathrm{gen}}_{\gamma}(H,S)\leq\frac{M^{2}\alpha}{2n}-\frac{\mathrm{Var}(\exp(\gamma\ell(H,Z)))}{2\gamma}\exp(-2\gamma M).$ (107)
Proof.
Let us consider the following decomposition,
$\begin{split}\overline{\mathrm{gen}}_{\gamma}(H,S)&=\mathrm{R}(H,P_{H}\otimes\mu)-\mathbb{E}_{P_{H,S}}[\frac{1}{n}\sum_{i=1}^{n}\ell(H,Z_{i})]\\ &\quad+\mathbb{E}_{P_{H,S}}[\frac{1}{n}\sum_{i=1}^{n}\ell(H,Z_{i})]-\overline{\mathrm{R}}_{\gamma}(H,P_{H,S}).\end{split}$ (108)
From Aminian et al. (2021a), for the Gibbs posterior we have
$\mathrm{R}(H,P_{H}\otimes\mu)-\mathbb{E}_{P_{H,S}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\ell(H,Z_{i})\Big{]}\leq\frac{\alpha M^{2}}{2n}.$
In addition, using Lemma C.3 for uniform distribution, we have
$\begin{split}&\frac{1}{n}\sum_{i=1}^{n}\frac{1}{\gamma}\log\big{(}\exp(\gamma\ell(H,Z_{i}))\big{)}-\frac{1}{\gamma}\log\Big{(}\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))\Big{)}\\ &\leq\frac{-\mathrm{Var}(\exp(\gamma\ell(H,Z)))}{2\gamma}\exp(-2\gamma M).\end{split}$
This completes the proof. ∎
From Proposition G.2, we can observe that for $\gamma>0$ the upper bound on the tilted generalization error under the Gibbs posterior is tighter than the upper bound on the generalization error of the Gibbs posterior under linear empirical risk given in Aminian et al. (2021a).
Furthermore, we can provide an upper bound on the absolute value of the expected tilted generalization error under the Gibbs posterior,
$\big{|}\overline{\mathrm{gen}}_{\gamma}(H,S)\big{|}\leq\frac{M^{2}\alpha}{2n}+\frac{\max(1,\exp(-2\gamma M))}{8\gamma}(1-\exp(\gamma M))^{2}.$ (109)
In (109), choosing $\gamma=O(1/n)$ we obtain a proof of a convergence rate of $O(1/n)$ for the upper bound on the absolute value of the expected tilted generalization error of the Gibbs posterior.
Appendix H Other Bounds
In this section, we provide upper bounds via Rademacher complexity, stability and PAC-Bayesian approaches. The results are based on the assumption of bounded loss functions (Assumption 3.1).
H.1 Rademacher Complexity
Inspired by the work (Bartlett and Mendelson, 2002), we provide an upper bound on the tilted generalization error via Rademacher complexity analysis. For this purpose, we need to define the Rademacher complexity.
As in Bartlett and Mendelson (2002), for a hypothesis set $\mathcal{H}$ of functions $h:\mathcal{X}\mapsto\mathcal{Y}$, the Rademacher complexity with respect to the dataset $S$ is
$\mathfrak{R}_{S}(\mathcal{H}):=\mathbb{E}_{S,\boldsymbol{\sigma}}\Big{[}\sup_{h\in\mathcal{H}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}h(X_{i})\Big{]}\,,$
where $\boldsymbol{\sigma}=\{\sigma_{i}\}_{i=1}^{n}$ are i.i.d Rademacher random variables; $\sigma_{i}\in\{-1,1\}$ and $\sigma_{i}=1$ or $\sigma_{i}=-1$ with probability 1/2, for $i\in[n]$. The empirical Rademacher complexity ^$\hat{\mathfrak{R}}_{S}(\mathcal{H})$ with respect to $S$ is defined by
^$\hat{\mathfrak{R}}_{S}(\mathcal{H}):=\mathbb{E}_{\boldsymbol{\sigma}}\Big{[}\sup_{h\in\mathcal{H}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}h(X_{i})\Big{]}\,.$ (110)
To provide an upper bound on the tilted generalization error, first, we apply the uniform bound, Lemma C.11, and Talagrand’s contraction lemma (Talagrand, 1996) in order to derive a high-probability upper bound on the tilted generalization error; we employ the notation (110).
Proposition H.1.
Given Assumptions 3.1 and assuming the loss function is $M_{\ell^{\prime}}$-Lipschitz-continuous in a binary classification problem, the tilted generalization error satisfies with probability at least $(1-\delta)$ that $\begin{split}\widehat{\mathrm{gen}}_{\gamma}(h,S)\leq 2\exp(|\gamma|M)M_{\ell^{\prime}}\hat{\mathfrak{R}}_{\mathrm{S}}(\mathcal{H})+\frac{3(\exp(|\gamma|M)-1)}{|\gamma|}\sqrt{\frac{\log(1/\delta)}{2n}}.\end{split}$
Proof.
Note that $\exp(\gamma M)\leq x\leq 1$ for $\gamma<0$ and $1\leq x\leq\exp(\gamma M)$ for $\gamma>0$. Therefore, we have the Lipschitz constant $\exp(-\gamma M)$ and $1$ for negative and positive $\gamma$, respectively. Similarly, for $\exp(\gamma x)$ and $0<x<M$, we have the Lipschitz constants $\gamma$ and $\gamma\exp(\gamma M)$, for $\gamma<0$ and $\gamma>0$, respectively. For $\gamma<0$, we have
[…]
where (a) holds due to the Lipschitzness of $\log(x)$ in a bounded interval, (b) holds due to the uniform bound Lemma C.11, (c) and (d) hold due to Talagrand’s contraction Lemma C.12).
Similarly, we can prove for $\gamma>0$, we have
$\begin{split}\widehat{\mathrm{gen}}_{\gamma}(h,S)\leq 2\exp(\gamma M)M_{\ell^{\prime}}\hat{\mathfrak{R}}_{\mathrm{S}}(\mathcal{H})+\frac{3(\exp(\gamma M)-1)}{\gamma}\sqrt{\frac{\log(1/\delta)}{2n}}.\end{split}$ (112)
Then, we obtain an upper bound on the generalization error by combining Proposition H.1, Massart’s lemma (Massart, 2000) and Lemma C.3.
Theorem H.2.
Under the same assumptions as in Proposition H.1, assuming a finite hypothesis space, the tilted generalization error satisfies with probability at least $(1-\delta)$ that $\begin{split}\mathrm{gen}_{\gamma}(h,S)\leq&\frac{\max(1,\exp(-2\gamma M))}{8\gamma}(\exp(\gamma M)-1)^{2}+2AM_{\ell^{\prime}}B\frac{\sqrt{2\log(\mathrm{card}(\mathcal{H}))}}{n}\\ &+\frac{3(A(\gamma)-1)}{|\gamma|}\sqrt{\frac{\log(1/\delta)}{2n}},\end{split}$ where $A(\gamma)=\exp(|\gamma|M)$ and $B^{2}=\max_{h\in\mathcal{H}}\Big{(}\sum_{i=1}^{n}h^{2}(z_{i})\Big{)}$.
Proof.
We consider the following decomposition for the Rademacher complexity,
$\mathrm{gen}_{\gamma}(h,S)=\mathrm{R}(h,\mu)-\mathrm{R}_{\gamma}(h,\mu^{\otimes n})+\mathrm{R}_{\gamma}(h,\mu^{\otimes n})-\widehat{\mathrm{R}}_{\gamma}(h,S),$
where $\mathrm{R}(h,\mu)-\mathrm{R}_{\gamma}(h,\mu^{\otimes n})$ can be bounded using Proposition 3.2. The second term can be bounded by using Proposition H.1 and Massart’s lemma (Lemma C.13). ∎
Similar to Remark 3.8, assuming $\gamma=O(1/\sqrt{n})$, we have the convergence rate of $O(1/\sqrt{n})$ for the tilted generalization error. For an infinite hypothesis space, covering number bounds can be applied to the empirical Rademacher complexity, see, e.g., (Kakade et al., 2008). We note that the VC-dimension and Rademacher complexity bounds are uniform bounds and are independent of the learning algorithms.
H.2 A Stability Bound
In this section, we also study the upper bound on the tilted generalization error from the stability perspective (Bousquet and Elisseeff, 2002b). In the stability approach, (Bousquet and Elisseeff, 2002b), the learning algorithm is a deterministic function of $S$.
For stability analysis, we define the replace-one sample dataset as
$S_{(i)}=\{Z_{1},\cdots,\tilde{Z}_{i},\cdots,Z_{n}\},$
where the sample $Z_{i}$ is replaced by an i.i.d. data sample $\tilde{Z}_{i}$ sampled from $\mu$. To distinguish the hypothesis in the stability approach from the uniform approaches, we consider $h_{s}:\mathcal{Z}^{n}\mapsto\mathcal{H}$ as the learning algorithm. In the stability approach, the hypothesis is a deterministic function $h_{s}(S)$ of the dataset. We are interested in providing an upper bound on the expected tilted generalization error $\mathbb{E}_{P_{S}}\big{[}\mathrm{gen}_{\gamma}(h_{s}(S),S)\big{]}$.
Theorem H.3.
Under Assumption 3.1, the following upper bound holds with probability at least $(1-\delta)$ under distribution $P_{S}$, $\begin{split}&\mathbb{E}_{P_{S}}\big{[}\mathrm{gen}_{\gamma}(h_{s}(S),S)\big{]}\\ &\leq\frac{(1-\exp(\gamma M))^{2}}{8\gamma}\Big{(}1+\exp(-2\gamma M)\Big{)}+\exp(|\gamma|M)\mathbb{E}_{P_{S},\tilde{Z}}[|\ell(h_{s}(S),\tilde{Z})-\ell(h_{s}(S_{(i)}),\tilde{Z})|].\end{split}$ (113)
Proof.
We use the following decomposition of the tilted generalization error;
$\begin{split}&\mathbb{E}_{P_{S}}\big{[}\mathrm{gen}_{\gamma}(h_{s}(S),S)\big{]}\\ &=\mathbb{E}_{P_{S}}\big{[}\mathrm{R}(h_{s}(S),\mu)-\frac{1}{\gamma}\log(\mathbb{E}_{P_{S},\mu}[\exp(\gamma\ell(h_{s}(S),\tilde{Z}))])\big{]}\\ &\quad+\mathbb{E}_{P_{S}}\bigg{[}\frac{1}{\gamma}\log(\mathbb{E}_{P_{S},\mu}[\exp(\gamma\ell(h_{s}(S),\tilde{Z}))])-\frac{1}{\gamma}\log\Big{(}\mathbb{E}_{P_{S}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(h_{s}(S),Z_{i}))\Big{]}\Big{)}\bigg{]}\\ &\quad+\mathbb{E}_{P_{S}}\bigg{[}\frac{1}{\gamma}\log\Big{(}\mathbb{E}_{P_{S}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(h_{s}(S),Z_{i}))\Big{]}\Big{)}-\widehat{\mathrm{R}}_{\gamma}(h_{s}(S),S)\bigg{]}.\end{split}$ (114)
Using Lemma C.3, we have
$\begin{split}&\mathbb{E}_{P_{S}}\big{[}\mathrm{R}(h_{s}(S),\mu)-\frac{1}{\gamma}\log(\mathbb{E}_{P_{S},\mu}[\exp(\gamma\ell(h_{s}(S),\tilde{Z}))])\big{]}\\ &\leq\frac{-\exp(-2\gamma M)}{2\gamma}\mathrm{Var}_{P_{S},\mu}(\exp(\gamma\ell(h_{s}(S),\tilde{Z}))),\end{split}$
and
$\begin{split}&\mathbb{E}_{P_{S}}\bigg{[}\frac{1}{\gamma}\log\Big{(}\mathbb{E}_{P_{S}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(h_{s}(S),Z_{i}))\Big{]}\Big{)}-\widehat{\mathrm{R}}_{\gamma}(h_{s}(S),S)\bigg{]}\\ &=\frac{1}{\gamma}\log\Big{(}\mathbb{E}_{P_{S}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(h_{s}(S),Z_{i}))\Big{]}\Big{)}-\mathbb{E}_{P_{S}}\bigg{[}\frac{1}{\gamma}\log\Big{(}\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(h_{s}(S),Z_{i}))\Big{)}\bigg{]}\\ &\leq\frac{1}{2\gamma}\mathrm{Var}\big{(}\exp(\gamma\ell(h_{s}(S),Z_{i}))\big{)}.\end{split}$
Using the Lipschitz property of the log and exponential functions on a closed interval, we have
$\begin{split}&\Big{|}\frac{1}{\gamma}\log(\mathbb{E}_{P_{S},\mu}[\exp(\gamma\ell(h_{s}(S),\tilde{Z}))])-\frac{1}{\gamma}\log\Big{(}\mathbb{E}_{P_{S}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(h_{s}(S),Z_{i}))\Big{]}\Big{)}\Big{|}\\ &=\Big{|}\frac{1}{\gamma}\log(\mathbb{E}_{P_{S},\mu}[\exp(\gamma\ell(h_{s}(S),\tilde{Z}))])-\frac{1}{\gamma}\log\Big{(}\mathbb{E}_{P_{S}}\Big{[}\exp(\gamma\ell(h_{s}(S),Z_{i}))\Big{]}\Big{)}\Big{|}\\ &\leq\exp(|\gamma|M)\mathbb{E}_{P_{S},\mu}[|\ell(h_{s}(S),\tilde{Z})-\ell(h_{s}(S_{(i)}),\tilde{Z})|].\end{split}$
Finally, we have
$\begin{split}&\mathbb{E}_{P_{S}}\big{[}\mathrm{gen}_{\gamma}(h_{s}(S),S)\big{]}\\ &\leq\frac{1}{2\gamma}\mathrm{Var}\big{(}\exp(\gamma\ell(h_{s}(S),Z_{i}))\big{)}-\frac{\exp(-2\gamma M)}{2\gamma}\mathrm{Var}_{P_{S},\mu}(\exp(\gamma\ell(h_{s}(S),\tilde{Z})))\\ &\qquad+\exp(|\gamma|M)\mathbb{E}_{P_{S},\mu}[|\ell(h_{s}(S),\tilde{Z})-\ell(h_{s}(S_{(i)}),\tilde{Z})|]\\ &\leq\frac{(1-\exp(\gamma M))^{2}}{8\gamma}\Big{(}1+\exp(-2\gamma M)\Big{)}+\exp(|\gamma|M)\mathbb{E}_{P_{S},\mu}[|\ell(h_{s}(S),\tilde{Z})-\ell(h_{s}(S_{(i)}),\tilde{Z})|].\end{split}$
We also consider the uniform stability as in Bousquet and Elisseeff (2002b).
Definition H.4 (Uniform Stability).
A learning algorithm is uniform $\beta$-stable with respect to the loss function if the following holds for all $S\in\mathcal{Z}^{n}$ and $\tilde{z}_{i}\in\mathcal{Z}$,
$\big{|}\ell(h_{s}(S),\tilde{z}_{i})-\ell(h_{s}(S_{(i)}),\tilde{z}_{i})\big{|}\leq\beta,\quad i\in[n].$
Remark H.5 (Uniform Stability).
Suppose that the learning algorithm is $\beta$-uniform stable with respect to a given loss function. Then, using Theorem H.3, we have
$\begin{split}&\mathbb{E}_{P_{S}}\big{[}\mathrm{gen}_{\gamma}(h_{s}(S),S)\big{]}\leq\frac{(1-\exp(\gamma M))^{2}}{8|\gamma|}\Big{(}1+\exp(-2\gamma M)\Big{)}+\exp(|\gamma|M)\beta.\end{split}$ (115)
Note that for a learning algorithm with uniform $\beta$-stability, where $\beta=O(1/n)$, then with $\gamma$ of order $O(1/n)$, we obtain a guarantee on the convergence rate of $O(1/n)$.
H.3 A PAC-Bayesian Bound
Inspired by previous works on PAC-Bayesian theory, see, e.g.,(Alquier, 2021; Catoni, 2003), we derive a high probability bound on the expectation of the tilted generalization error with respect to the posterior distribution over the hypothesis space.
In the PAC-Bayesian approach, we fix a probability distribution over the hypothesis (parameter) space as prior distribution, denoted as $Q_{h}$. Then, we are interested in the generalization performance under a data-dependent distribution over the hypothesis space, known as posterior distribution, denoted as $\rho_{h}$.
Theorem H.6.
Under Assumption 3.1, the following upper bound holds on the conditional expected tilted generalization error with probability at least $(1-\delta)$ under the distribution $P_{S}$; for any $\eta>0$, $\begin{split}&|\mathbb{E}_{\rho_{h}}[\mathrm{gen}_{\gamma}(H,S)]|\leq\frac{\max(1,\exp(-2\gamma M))(1-\exp(\gamma M))^{2}}{8|\gamma|}\\ &\qquad+\frac{L\eta A^{2}}{8n}+\frac{L(\mathrm{KL}(\rho_{h}\|Q_{h})+\log(1/\delta))}{\eta},\end{split}$ (116) where $Q_{h}$ and $\rho_{h}$ are prior and posterior distributions over the hypothesis space, respectively.
Proof.
We use the following decomposition of the generalization error,
$\begin{split}\mathbb{E}_{\rho_{h}}[\mathrm{gen}_{\gamma}(H,S)]=\mathbb{E}_{\rho_{h}}[\mathrm{R}(H,\mu)-\mathrm{R}_{\gamma}(H,\mu)+\mathrm{R}_{\gamma}(H,\mu)-\widehat{\mathrm{R}}_{\gamma}(H,S)].\end{split}$
The term $\mathbb{E}_{\rho_{h}}[\mathrm{R}(H,\mu)-\mathrm{R}_{\gamma}(H,\mu)]$ can be bounded using Lemma C.3. The second term $\mathrm{R}_{\gamma}(H,\mu)-\widehat{\mathrm{R}}_{\gamma}(H,S)$ can be bounded using the Lipschitz property of the log function and Catoni’s bound (Catoni, 2003). ∎
Remark H.7.
Choosing $\eta$ and $\gamma$ such that $\eta^{-1}\asymp 1/\sqrt{n}$ and $\gamma=O(1/\sqrt{n})$ results in a theoretical guarantee on the convergence rate of $O(1/\sqrt{n})$.